### v1

# Stability via resampling: statistical problems beyond the real line

Abstract
Model averaging techniques based on resampling methods (such as bootstrapping or subsampling) have been utilized across many areas of statistics, often with the explicit goal of promoting stability in the resulting output. We provide a general, finite-sample theoretical result guaranteeing the stability of bagging when applied to algorithms that return outputs in a general space, so that the output is not necessarily a real-valued—for example, an algorithm that estimates a vector of weights or a density function. We empirically assess the stability of bagging on synthetic and real-world data for a range of problem settings, including causal inference, nonparametric regression, and Bayesian model selection.
1 Introduction
Statistical stability holds when our conclusions, estimates, fitted models, predictions, or decisions are insensitive to small perturbations of data. Stability is a pillar of modern data science (Yu and Kumbier, 2020). Nonparametric tools for statistical inference—e.g. the bootstrap, subsampling, the Jackknife—perturb data to assess variability and can thus be viewed as aspects of stability analysis (Yu, 2013). While assessing stability is in general statistically intractable without placing assumptions on the data (Kim and Barber, 2023), these resampling-based approaches can also be used to reduce the issue of instability: for instance, Breiman (1996a, b) originally proposed bagging—which applies bootstrap resampling to construct an ensemble of machine learning models—as an off-the-shelf stabilizer for regression and classification. In this paper, we show how bagging can be used to stabilize statistical methods far beyond its origins in supervised learning problems.
1.1 Problem settings
We begin by highlighting some specific statistical problems where provably stable methods are desirable.
Problem setting 1 (background): the supervised learning setting
In the supervised learning setting, we are given a labeled data set $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ and are tasked with constructing a regression function $\hat{f}$ to predict $Y$ given $X$ for new data points. Many modern methods from the recent machine learning literature offer powerful predictive performance, but suffer from high sensitivity to the training data, where a prediction $\hat{f}(X)$ (at a new test point $X$) may vary immensely if the training data is perturbed slightly. Interestingly, both high accuracy and poor stability can often be a consequence of working with highly overparameterized models.
Suppose that the response $Y$ is real-valued, so that our regression functions take values in $\mathbb{R}$, i.e., $\hat{f}:\mathcal{X}\to\mathbb{R}$. Recent work by Soloff et al. (2023) establishes that bagging the predictive model—that is, re-fitting $\hat{f}$ on bootstrapped training sets, and then returning the averaged prediction—yields an assumption-free guarantee on the stability of the prediction $\hat{f}(X)\in\mathbb{R}$.
While Soloff et al. (2023)’s work is presented as a result for supervised learning, it can be viewed more generally as a guarantee of stability for any algorithm that is trained on data and returns a real-valued output. But even with this generalization, this result only covers a very specific statistical scenario. For many statistical problems, we might be interested in the stability of a procedure that returns a more complex object, such as a function or a vector of weights, rather than a real-valued output.
Problem setting 2: reproducible Bayesian inference
In Bayesian data analysis, given a data set $\mathcal{D}$ containing $n$ observations drawn from some distribution $p_{\theta}$, we are interested in the posterior distribution of the parameter $\theta$. For instance, $\theta$ might represent a particular model from a finite set of possibilities, in which case the posterior distribution on $\theta$ is simply a vector of weights on the finite set of possibilities. Alternatively, $\theta$ might take values in $\mathbb{R}^{d}$ (e.g., coefficients in some regression model), in which case the posterior distribution is a distribution over $\mathbb{R}^{d}$—a more complex object.
Bayesian inference may sometimes lead to answers that are highly sensitive with respect to perturbations of the data. In particular, this can occur in settings where the model is misspecified (but nonetheless, is a reasonably good approximation to the observed data, so that computing the posterior is still a reasonable goal). To alleviate this instability, the BayesBag approach (Bühlmann, 2014) averages posterior distributions over different resamplings of the data $\mathcal{D}$. Huggins and Miller (2023) analyzes the accuracy and stability properties of the BayesBag approach for the specific problem of Bayesian model selection, where $\theta$ lies in a finite set of models $\mathfrak{M}=\{\mathfrak{m}_{1},\dots,\mathfrak{m}_{K}\}$ (and so the posterior on $\theta$ is given by a vector of weights of length $K$, i.e., with the $k$th weight indicating the posterior probability $\mathbb{P}\{\theta=\mathfrak{m}_{k}\mid\mathcal{D}\}$). Their work establishes BayesBag as an empirically effective method with asymptotic guarantees, but has not yielded finite-sample guarantees even in the finite model setting. Of course, we can also ask this question even more generally: will BayesBag lead to greater stability in a general setting, where $\theta$ may lie in an arbitrary space (e.g., if the posterior is a distribution on $[0,1]$, or on $\mathbb{R}^{d}$, etc.)?
Problem setting 3: synthetic controls for causal inference
The synthetic control method (SCM) is a popular approach for estimating causal effects with panel data (Abadie and Gardeazabal, 2003; Abadie et al., 2010, 2015; Abadie, 2021). The method is particularly useful in settings with a single treated unit and a relatively small number of control units. A synthetic control is a weighted combination of control units, where the weights are chosen to match the treated unit’s pre-treatment outcomes and other covariates. SCM estimates the counterfactual—what outcomes would we expect for the treated unit had the policy not been adopted?—using the synthetic control’s post-treatment outcomes. For example, Abadie et al. (2015) applies SCM to the 1990 German reunification and its impact on per capita GDP in West Germany. Per capita GDP was increasing prior to reunification and continued to increase in subsequent years, for this was a period of sustained growth across industrialized countries. To assess the counterfactual, a ‘synthetic West Germany’ is constructed as follows: we find the convex combination of other industrialized countries such that the pre-1990 economic trend most closely resembles that of West Germany. Abadie et al. (2015) finds that West Germany’s economy did not grow as much as it would have without reunification.
Often SCM’s weights are of direct interest, as they say something about which peer countries are most relevant in this counterfactual assessment. However, in order to reliably interpret the weights, we should first ensure they are stable (Murdoch et al., 2019). For instance, if we remove Denmark from the analysis (a country which receives zero weight in the synthetic West Germany) then there is no longer a positive weight on the United States, which previously received $22\%$ of the weight. This sensitivity appears to betray our intuitive interpretation of the meaning of SCM’s weights. Can a resampling-type approach to SCM alleviate these issues?
1.2 Our contributions
Algorithmic stability is desirable for statistical methods far beyond its origins in statistical learning theory (Devroye and Wagner, 1979a, b; Bousquet and Elisseeff, 2002). Broadly, stability is considered a ‘prerequisite for trustworthy interpretations’ (Murdoch et al., 2019). The goal of this paper is to establish quantitative guarantees on the stability of bagging as a general-purpose tool across a wide range of problems where the output—the output being estimated from the data—might not be real-valued and might even be infinite-dimensional. Table 1 lists a range of problem settings where our framework can be directly applied (including the settings discussed above, namely, supervised learning, Bayesian inference, and synthetic controls). Formally, our results will apply to any algorithm, randomized or deterministic, with outputs in bounded subset of a separable Hilbert or Banach space—we give details below. We highlight one particularly important finding: our main results will show that for a Hilbert space, there is no cost to high dimensionality in terms of the stability guarantees that can be obtained (in contrast, for the more general setting of a Banach space, this is no longer the case).
Outline
The remainder of this paper is organized as follows. In Section 2, we define a framework for algorithmic stability in any metric space, and then define bagging for an arbitrary algorithm in Section 3. Our main results on the stability of bagging are presented in Section 4 and Section 5. In Section 6, we take a closer look at the implications of our results on a range of specific settings appearing in Table 1, and discuss some consequences of our stability guarantee. We conclude with a discussion in Section 7. All proofs are deferred to the Appendix.
Applications Output Space $\mathcal{W}$ Metric $\mathsf{d}$ on $\mathcal{W}$
Supervised learning Predicted value range $[a,b]\subset\mathbb{R}$ Absolute value ||$\cdot$|| ⋅ |
Bayesian model selection
Synthetic controls
Probability simplex $\Delta_{K-1}$
Euclidean dist. $\|\cdot\|_{2}$ or
Total variation dist. $\mathsf{d}_{\mathsf{TV}}$
Constrained Lasso ($q=1$)
Constrained Ridge ($q=2$)
$\ell_{q}$-ball $\mathbb{B}_{q}^{d}(R)$ $\ell_{q}$-norm $\|\cdot\|_{q}$
Density estimation $L$-Lipschitz densities on $[0,1]$ Total variation dist. $\mathsf{d}_{\mathsf{TV}}$
Nonparametric regression Sobolev ball of functions $\mathbb{R}^{d}\to\mathbb{R}$ Sobolev norm $\|\cdot\|_{s,2}$
Table 1: Example problem settings for our stability framework. The first three examples correspond to Problem Settings 1-3 described in Section 1.
2 Framework: algorithmic stability in a metric space
We define a randomized algorithm $\mathcal{A}$ as function that takes in a finite sequence of data $\mathcal{D}=(Z_{i})_{i\in[n]}$, as well as some independent randomness $\xi\sim\text{Uniform}([0,1])$, and returns an output $\mathcal{A}(\mathcal{D};\xi)$. The term $\xi$, which we refer to as the random seed, determines all randomness used by the algorithm—for instance, a random initialization or a random split of the data into batches. The algorithm $\mathcal{A}$ is deterministic if it does not depend on the seed $\xi$.
This notion of a (deterministic or randomized) algorithm covers many disparate problems in statistical learning, such as the problem settings introduced in Section 1.1.
  • Consider a supervised learning setting (Problem Setting 1) with data points $Z_{i}=(X_{i},Y_{i})\in\mathcal{X}\times\mathbb{R}$. If we are interested in predicting $Y$ given a new feature $X=x$, our algorithm $\mathcal{A}$ can return $\hat{f}(x)\in\mathbb{R}$, the predicted value at $X=x$, as its output. (Alternatively, if we are interested in building a model for $Y$ given $X$, our algorithm $\mathcal{A}$ can return an entire fitted regression function $\hat{f}$—a more complex output.)
  • In Bayesian inference (Problem Setting 2), we observe data points $Z_{1},\dots,Z_{n}$ drawn from some distribution $p_{\theta}$, where $\theta\in\Theta$ is the unknown parameter. To perform Bayesian inference on $\theta$, our algorithm $\mathcal{A}$ would then return the posterior distribution over $\theta\in\Theta$.
  • For the problem of synthetic controls (Problem Setting 3), our data points $Z_{i}$ are the feature vectors for each control unit $i=1,\dots,n$, and the algorithm $\mathcal{A}$ returns a weight vector of dimension $n$, specifying the convex combination of control units that define the synthetic control.
In some settings it may be common to use a randomized, rather than deterministic, algorithm. For example, in Bayesian inference, in a high-dimensional setting we might choose to approximate the posterior through (random) sampling rather than computing the density explicitly.
Formally, we write
$\mathcal{A}:\bigcup_{n\geq 0}\mathcal{Z}^{n}\times[0,1]\to\mathcal{W},$
where the output takes values in some space $\mathcal{W}$ equipped with a metric $\mathsf{d}$—see Table 1 for examples. From this point on, we write $\hat{w}=\mathcal{A}(\mathcal{D};\xi)$ to denote the output, regardless of the nature of the output (e.g., a finite-dimensional vector of weights for the synthetic control problem, or a posterior density function for the Bayesian inference setting, etc.). [^1] We assume throughout that $t\mapsto\mathcal{A}((Z_{i})_{i\in[n]},t)$, which is a function from $[0,1]$ to $\mathcal{W}$, is measurable for any fixed $(Z_{i})_{i\in[n]}\in\mathcal{Z}^{n}$. Measurability is defined with respect to the Borel $\sigma$-algebra on $(\mathcal{W},\mathsf{d})$.
2.1 Defining stability
There are many related definitions of algorithmic stability in the literature. We focus on stability of the algorithm’s output with respect to dropping a data point: will the output $\hat{w}$ of the algorithm be similar if one data point is removed at random from the training data set?
Definition 1.
An algorithm $\mathcal{A}$ has mean-square stability $\beta^{2}$ with respect to a metric $\mathsf{d}$ if, for all data sets $\mathcal{D}\in\mathcal{Z}^{n}$ of size $n$,
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\xi}\left[\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\right]\leq\beta^{2}.$
where $\hat{w}=\mathcal{A}(\mathcal{D};\xi)$ is the output of the algorithm run on the full data set $\mathcal{D}$, $\hat{w}^{\setminus i}=\mathcal{A}(\mathcal{D}^{\setminus i};\xi)$ is the output of the algorithm run on the leave-one-out data set $\mathcal{D}^{\setminus i}=(Z_{j})_{j\neq i}$ and the expectation averages over $\xi\sim\text{Uniform}[0,1]$.
This first definition captures the average perturbation in $\hat{w}$ when one data point is removed at random from some data set $\mathcal{D}\in\mathcal{Z}^{n}$—that is, $\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)$, averaged over which data point $i$ gets removed. An alternative way to define stability is to control the tails of these perturbations $\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)$, ensuring that not too many of them can be large.
Definition 2.
An algorithm $\mathcal{A}$ has tail stability $(\varepsilon,\delta)$ with respect to a metric $\mathsf{d}$ if, for all data sets $\mathcal{D}\in\mathcal{Z}^{n}$ of size $n$,
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}(\hat{w},\hat{w}^{\setminus i})>\varepsilon\right\}\leq\delta,$
where $\hat{w}=\mathcal{A}(\mathcal{D};\xi)$ and $\hat{w}^{\setminus i}=\mathcal{A}(\mathcal{D}^{\setminus i};\xi)$ as in Definition 1.
The terminology ‘$\mathcal{A}$ has mean-square stability $\beta^{2}$’ in Definition 1, and ‘$\mathcal{A}$ has tail stability $(\varepsilon,\delta)$’ in Definition 2, suppresses the dependence on the sample size $n$. Since we perform a non-asymptotic stability analysis, we treat $n$ as a fixed positive integer throughout.
These two notions of stability are clearly very similar in flavor, and the following simple result shows that each implies the other.
Proposition 3.
If $\mathcal{A}$ has mean-square stability $\beta^{2}$ with respect to $\mathsf{d}$, then $\mathcal{A}$ has tail stability $(\varepsilon,\delta)$ with respect to $\mathsf{d}$, for any $\varepsilon,\delta\geq 0$ satisfying $\delta\varepsilon^{2}\geq\beta^{2}$. Conversely, if $\mathcal{A}$ has tail stability $(\varepsilon,\delta)$ with respect to $\mathsf{d}$, then $\mathcal{A}$ has mean-square stability $\beta^{2}=\varepsilon^{2}+\delta\,\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right)$ with respect to $\mathsf{d}$.
Proof.
The first claim follows by Markov’s inequality, since
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)>\varepsilon\right\}=\mathbb{E}_{\xi}\left[\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)>\varepsilon^{2}\right\}\right]\\ \leq\varepsilon^{-2}\mathbb{E}_{\xi}\left[\frac{1}{n}\sum_{i=1}^{n}\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\right]\leq\varepsilon^{-2}\beta^{2}.$
For the second claim, since $\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\leq\varepsilon^{2}+\mathbf{1}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)>\varepsilon\right\}\cdot\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right)$, we have
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\xi}\,\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)$ $\leq\varepsilon^{2}+\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)>\varepsilon\right\}\cdot\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right)$
$\leq\varepsilon^{2}+\delta\cdot\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right),$
completing the proof. ∎
2.2 Related notions of stability
Our definitions of stability are stronger than many common definitions in the literature. For instance, we require the algorithm to be stable for any fixed data set $\mathcal{D}$ to certify the stability of the method without looking at the specific data set. Definition 2 implies in particular that the algorithm is stable when the data $(Z_{i})_{i\in[n]}$ are i.i.d. from some distribution $P$, but our definition does not require any stochastic model.
In many cases, practitioners care specifically about the stability of the algorithm’s output $\hat{w}$, e.g., if we want to interpret the values of some learned weights, cluster assignments, etc. However, in other cases, we may be more concerned with the stability of some downstream decisions or outcomes based on an analysis. For instance, in supervised learning (Problem Setting 1), we may desire stability of a loss incurred by the output $\hat{w}$—e.g., a predictive error $\ell(\hat{w},(x,y))=|y-\hat{w}(x)|$. These types of loss-based stability properties can often be inferred from stability of the output $\hat{w}$ (see Soloff et al. (2023) for further discussion). In this paper, we only consider the stability of $\hat{w}$ itself.
3 Constructing a bagged algorithm via resampling
In this section, we formally define the construction of a bagged algorithm.
Stability is about (in)sensitivity to perturbed data sets. Resampling methods deliberately perturb the data, so they are naturally a useful tool for stabilizing algorithms: given a data set $\mathcal{D}\in\mathcal{Z}^{n}$ containing $n$ data points, we might retrain our algorithm repeatedly on resampled data sets, and then average the resulting outputs. There are several common schemes for constructing these resampled data sets. Here, we mention two of the most common methods:
  • Bootstrapping constructs each resampled data set by sampling $m$ indices $i_{1},\dots,i_{m}$ uniformly with replacement from $[n]$ (the term “bootstrap” commonly refers to this method with the choice $m=n$, so that the resampled data sets are of the same size as the original $\mathcal{D}$).
  • Alternatively, we may instead sample data points without replacement, so that each contains a uniformly drawn subset containing $m$ of the original $n$ data points (e.g., with $m=n/2$). This type of approach is known as subbagging.
We now define some notation that unifies these two variants (as well as allowing for other resampling schemes, as we discuss below).
Define the set of finite sequences of indices in $[n]$:
$\textnormal{seq}_{[n]}:=\{(i_{1},\ldots,i_{k}):k\geq 0,\,i_{1},\ldots,i_{k}\in[n]\}.$
A sequence $r=(i_{1},\dots,i_{m})\in\textnormal{seq}_{[n]}$ is commonly referred to as a bag in the context of data set resampling. For any $r\in\textnormal{seq}_{[n]}$, define $\mathcal{D}^{r}=(Z_{i})_{i\in r}\in\mathcal{Z}^{m}$, the corresponding data set. Note that if the bag $r$ contains repeated indices (i.e., $i_{k}=i_{\ell}$ for some $k\neq\ell$), then the same data point from the original data set $\mathcal{D}$ will appear multiple times in $\mathcal{D}^{r}$.
We are now ready to define the bagged version of any base algorithm $\mathcal{A}$, using a particular resampling distribution $\mathcal{Q}_{n}$. [^2] While the term “bagging” is sometimes used as a synonym for bootstrap averaging (i.e., averaging outputs after sampling from the data set without replacement) in the literature, in this paper we will use “bagging” more broadly, allowing for any resampling distribution $\mathcal{Q}_{n}$. From this point on, we assume that $\mathcal{W}$ is a closed and compact subset of a separable Banach space—this technical condition ensures that taking an average or an expected value of outputs in $\mathcal{W}$ is well-defined.
Definition 4 (Bagging for an algorithm $\mathcal{A}$).
Given a data set $\mathcal{D}\in\mathcal{Z}^{n}$ and a distribution $\mathcal{Q}_{n}$ on $\textnormal{seq}_{[n]}$, return the output
$\hat{w}=\frac{1}{B}\sum_{b=1}^{B}\mathcal{A}(\mathcal{D}^{r_{b}};\xi_{b}),$
where $(r_{b},\xi_{b})\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}\mathcal{Q}_{n}\times\textnormal{Uniform}[0,1]$.
For example, bootstrapping corresponds to choosing $\mathcal{Q}_{n}$ as the uniform distribution over all sequences of length $m$, while subbagging takes $\mathcal{Q}_{n}$ as the uniform distribution over sequences of $m$ distinct elements.
For our theoretical analysis, it will be useful to consider taking $B\to\infty$, so as to “derandomize” the procedure.
Definition 5 (Derandomized bagging for an algorithm $\mathcal{A}$).
Given a data set $\mathcal{D}\in\mathcal{Z}^{n}$ and a distribution $\mathcal{Q}_{n}$ on $\textnormal{seq}_{[n]}$, return the output
$\hat{w}=\mathbb{E}_{r,\xi}[\mathcal{A}(\mathcal{D}^{r};\xi)],$
where the expected value is computed with respect to $(r,\xi)\sim\mathcal{Q}_{n}\times\text{Uniform}[0,1]$.
3.1 A general framework for bagging
The two variants discussed above, bagging and subbagging, are special cases of a broader framework, allowing for more flexibility in the choice of the resampling distribution $\mathcal{Q}_{n}$. (See Soloff et al. (2023) for a more complete discussion of this framework, and for examples of resampling methods beyond bagging and subbagging.) Here we formalize the conditions that $\mathcal{Q}_{n}$ needs to satisfy, in order for our main results to be applied.
Assumption 1.
Fix a sample size $n\geq 1$. The following conditions must be satisfied by the resampling distribution $\mathcal{Q}_{n}$.
  • Symmetry: for all $m\geq 1$, $i_{1},\ldots,i_{m}\in[n]$, and permutations $\sigma\in\mathcal{S}_{n}$,
    $\mathcal{Q}_{n}\{(i_{1},\ldots,i_{m})\}=\mathcal{Q}_{n}\{(\sigma(i_{1}),\ldots,\sigma(i_{m}))\}.$
  • Nontrivial subsampling: it holds that $p_{\mathcal{Q}_{n}}<1$, where we define
    $p_{\mathcal{Q}_{n}}=\mathbb{E}_{r\sim\mathcal{Q}_{n}}\left[\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{i\in r\}\right],$
    the expected fraction of data points appearing in a bag $r$ sampled from $\mathcal{Q}_{n}$.
  • Nonpositive covariance:
    $\textnormal{Cov}_{r\sim\mathcal{Q}_{n}}(\mathbf{1}_{i\in r},\mathbf{1}_{j\in r})\leq 0\textnormal{ for any $i\neq j\in[n]$}.$
  • Compatibility with the leave-one-out distribution: the corresponding distribution $\mathcal{Q}_{n-1}$ on $\textnormal{seq}_{[n-1]}$ must equal the distribution of $r\sim\mathcal{Q}_{n}$ conditional on the event $n\not\in r$.
Before proceeding, we comment on the last part of the assumption: why does $\mathcal{Q}_{n-1}$ appear in our assumptions on $\mathcal{Q}_{n}$? To study the stability of the bagged algorithm (Definition 4), it is not sufficient to consider only $\mathcal{Q}_{n}$, the distribution on $\textnormal{seq}_{[n]}$—when we compare to the result of the algorithm on the leave-one-out data set $\mathcal{D}^{\setminus i}$, we now have sample size $n-1$ (instead of $n$), and consequently we need to specify a distribution $\mathcal{Q}_{n-1}$ on $\textnormal{seq}_{[n-1]}$, as well.
The four conditions of 2 are all satisfied by the standard resampling schemes described above, with $p_{\mathcal{Q}_{n}}=1-\left(1-\frac{1}{n}\right)^{m}$ for bagging (sampling $m$ out of $n$ points uniformly with replacement), and $p_{\mathcal{Q}_{n}}=\frac{m}{n}$ for subbagging (sampling $m$ out of $n$ points uniformly without replacement).
4 Main results for a Hilbert space
We are now ready to present our stability guarantees for bagging. Our first main result establishes that derandomized bagging automatically provides mean-square stability for any algorithm with outputs in a bounded subset of a Hilbert space. In this section, we work with stability with respect to the Hilbert norm $\|\cdot\|_{\mathcal{H}}$. For instance,
  • If $\mathcal{W}\subset\mathbb{R}^{d}$, then the Hilbert norm may be chosen to be the Euclidean norm $\|\cdot\|_{2}$, or more generally, a Mahalanobis distance $\|w\|_{\mathcal{H}}=\sqrt{w^{\top}Sw}$ for a fixed positive definite matrix $S$.
  • If $\mathcal{W}\subset L_{2}(\mathbb{R}^{d})$, the space of square-integrable functions $\mathbb{R}^{d}\to\mathbb{R}$, then the Hilbert norm is given by the $L_{2}$ norm on functions, $\|w\|_{\mathcal{H}}=(\int_{\mathbb{R}^{d}}w(x)^{2}\;\mathsf{d}x)^{1/2}$. Alternatively, we can take a Sobolev space of functions $\mathbb{R}^{d}\to\mathbb{R}$ with smoothness parameter $s$ with the Hilbert norm being the Sobolev norm $\|\cdot\|_{s,2}$ in (3); see Experiment 3 in Section 6.1 for details.
Theorem 6.
Suppose $\mathcal{A}$ takes values in a convex and closed subset $\mathcal{W}\subset\mathcal{H}$, where $\mathcal{H}$ is a separable Hilbert space, and define
$\mathsf{rad}_{\mathcal{H}}(\mathcal{W})=\inf_{w\in\mathcal{W}}\sup_{w^{\prime}\in\mathcal{W}}\|w-w^{\prime}\|_{\mathcal{H}}$
as the radius of the set $\mathcal{W}$. Fix a resampling distribution $\mathcal{Q}_{n}$ on $\textnormal{seq}_{[n]}$, which satisfies 1. Then derandomized bagging, run with base algorithm $\mathcal{A}$ and resampling distribution $\mathcal{Q}_{n}$, has mean-square stability $\beta^{2}$ with respect to $\|\cdot\|_{\mathcal{H}}$, with
$\beta^{2}:=\frac{\mathsf{rad}^{2}_{\mathcal{H}}(\mathcal{W})}{n-1}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}.$ (1)
Combining this result with 3 immediately yields a guarantee on $(\varepsilon,\delta)$-stability.
Corollary 7.
Under the notation and assumptions of Theorem 6, derandomized bagging has tail stability $(\varepsilon,\delta)$ with respect to $\|\cdot\|_{\mathcal{H}}$ for any $\varepsilon,\delta\geq 0$ satisfying
$\delta\varepsilon^{2}\geq\frac{\mathsf{rad}_{\mathcal{H}}^{2}(\mathcal{W})}{n-1}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}.$ (2)
Corollary 7 generalizes the main result of Soloff et al. (2023), which focuses on the supervised learning setting where the data points $Z_{i}=(X_{i},Y_{i})$ contain features $X_{i}\in\mathcal{X}$ and a response $Y_{i}\in\mathbb{R}$, and the output of the algorithm $\mathcal{A}$ is a predicted value $\hat{f}(x)$, at a test point $x\in\mathcal{X}$—i.e., in that work, the output space $\mathcal{H}=\mathbb{R}$.
Comparing to prior bounds
The main result of (Soloff et al., 2023, Theorem 8), ensuring the stability of algorithms that return a real-valued output, can be recovered as a special case of Corollary 7 above, by taking $\mathcal{W}=[0,1]\subseteq\mathcal{H}=\mathbb{R}$ (note that we then have $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})=\frac{1}{2}$). But remarkably, the result of Theorem 6 shows that there is no price to pay for high-dimensionality (or even for infinite dimensionality): the stability guarantee is exactly the same for any Hilbert space, regardless of dimension, as long as the radius of the output space $\mathcal{W}$ is bounded.
4.1 Results for a finite number of bags
The results above apply to the derandomized bagged version of $\mathcal{A}$, as constructed in Definition 5. In practice, of course, it is not feasible to calculate this output $\hat{w}$ since we would need to average over all possible bags $r$, so we would instead use the bagged estimator $\hat{w}$ constructed in Definition 4 for some finite number of bags $B$. Since we are simply taking a Monte Carlo approximation to the derandomized bagged estimator of Definition 5, we would expect to incur error on the order of $\frac{1}{\sqrt{B}}$. To formalize this, we first state a generalization of the Azuma-Hoeffding inequality:
Theorem 8 ((Hayes, 2005)).
Let $\hat{w}_{1},\ldots,\hat{w}_{B}\in\mathcal{H}$ be independent random variables in a Hilbert space $\mathcal{H}$ such that $\|\hat{w}_{b}-\mathbb{E}\hat{w}_{b}\|_{\mathcal{H}}\leq C$ almost surely for $b=1,\ldots,B$. Then for any $\delta>0$,
$\mathbb{P}\left\{\bigg{\|}\frac{1}{B}\sum_{b=1}^{B}(\hat{w}_{b}-\mathbb{E}\hat{w}_{b})\bigg{\|}_{\mathcal{H}}\leq C\sqrt{\frac{1}{2B}\log\left(\frac{2e^{2}}{\delta}\right)}\right\}\geq 1-\delta.$
In particular,
$\mathbb{E}\left[\bigg{\|}\frac{1}{B}\sum_{b=1}^{B}(\hat{w}_{b}-\mathbb{E}\hat{w}_{b})\bigg{\|}_{\mathcal{H}}^{2}\right]\leq\frac{C^{2}e^{2}}{B}.$
Note that, aside from the additional factor $e^{2}$, this result is the same as what we would obtain for the case $\mathcal{H}=\mathbb{R}$. Combining this result with Theorem 6 yields the following stability guarantee for bagged algorithms with a finite number of bags $B$.
Corollary 9.
Under the notation and assumptions of Theorem 6, let $B\geq 1$ be fixed. Then bagging, run with base algorithm $\mathcal{A}$ and resampling distribution $\mathcal{Q}_{n}$ and with $B$ bags, has mean-square stability $\beta^{2}$ with respect to $\|\cdot\|_{\mathcal{H}}$, with
$\beta^{2}:=\mathsf{rad}^{2}_{\mathcal{H}}(\mathcal{W})\left(\frac{1}{n-1}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{16e^{2}}{B}\right).$
Of course, tail stability follows as an immediate consequence of 3.
5 Extension: results outside the Hilbert space setting
[…]
As a motivating example, consider a setting where $\mathcal{A}$ returns an output $\hat{w}$ that specifies a distribution—for instance, this may arise in Bayesian inference (Problem Setting 2 from Section 1.1), where $\mathcal{A}$ returns a posterior distribution over some parameter of interest. In this type of setting, since $\hat{w}$ represents a distribution, it is natural to study its stability with respect to $\mathsf{d}_{\mathsf{TV}}$, the total variation distance. While $\mathsf{d}_{\mathsf{TV}}$ defines a valid norm, this is not a Hilbert norm and so the theoretical guarantees of Section 4 cannot be directly applied.
[…]
5.1 Theoretical guarantee
Unlike the special case of a Hilbert space, in this more general setting we need an additional assumption to ensure that stability will hold.
Assumption 2.
There exists a constant $R>0$ such that $\sup_{w,w^{\prime}\in\mathcal{W}}\|w-w^{\prime}\|\leq R$, and moreover, there exist some fixed elements $v_{1},\dots,v_{K}\in\mathcal{B}$ with unit norm, $\|v_{1}\|=\dots=\|v_{K}\|=1$, and a constant $\rho>0$ such that for any $w,w^{\prime}\in\mathcal{W}$, the difference $w-w^{\prime}$ can be approximated by a linear combination of the $v_{j}$’s,
$\Big{\|}(w-w^{\prime})-\sum_{i=1}^{K}t_{i}v_{i}\Big{\|}\leq\rho$
for some $t=(t_{1},\dots,t_{K})\in\mathbb{R}^{K}$ with $\|t\|_{1}\leq R$.
For instance, returning to our earlier example, if $\mathcal{W}=\Delta_{K-1}$ (and the norm is given by $\mathsf{d}_{\mathsf{TV}}$) then we can take $v_{j}$ to be the $j$th canonical basis vector, and then 2 holds, with this same value of $K$, and with $R=1$ and $\rho=0$. Here $R$ is playing a role analogous to the radius, $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})$, in the Hilbert space setting.
Theorem 10.
[…]
$\beta^{2}:=1.6\left[R\sqrt{\frac{2(1+\eta_{n})\log(2K)}{n}\left(\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{4\eta_{n}}{(1-p_{\mathcal{Q}_{n}})^{2}}\right)}+R\sqrt{\frac{1}{n}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}+2\rho\cdot p_{\mathcal{Q}_{n}}\right]^{2},$
where $\eta_{n}=\frac{H_{n}-1}{n-H_{n}}\asymp\frac{\log n}{n}$, for $H_{n}=1+\frac{1}{2}+\dots+\frac{1}{n}$ denoting the $n$th harmonic number.
As in the Hilbert space setting, we can also extend this theorem to verify tail stability and to prove a stability result for bagging with a finite $B$—see Appendix B for details.
5.2 Examples
Now we return to Problem Setting 2, as discussed in the motivation for the extension to Banach spaces, where $\mathcal{A}$ returns a distribution (e.g., a posterior), to see what the new theorem implies in this setting. To make this more concrete, we can consider two specific scenarios: a discrete setting and a continuous setting. We will see how, in each scenario, the new result of Theorem 10 gives a much stronger guarantee than what can be derived from the results of Section 4. For intuition, in this discussion, we treat $\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}$ as a constant.
Problem Setting 2(a): the discrete case
First, consider a finite setting, where $\mathcal{A}$ returns a distribution on $K$ many values—for instance, if the output of $\mathcal{A}$ specifies a Bayesian posterior over $K$ many possible models. In this scenario, $\hat{w}$ takes values in the simplex $\Delta_{K-1}\subset\mathbb{R}^{K}$, and so the total variation distance is given by $\mathsf{d}_{\mathsf{TV}}(w,w^{\prime})=\frac{1}{2}\|w-w^{\prime}\|_{1}$.
Of course, since $\mathbb{R}^{K}$ is itself a Hilbert space (with norm $\|\cdot\|_{\mathcal{H}}=\|\cdot\|_{2}$), we can apply our results from Section 4—but this will not give a very useful bound for the stability of derandomized bagging. Specifically, Theorem 6 tells us that mean-square stability with respect to $\|\cdot\|_{2}$ holds, for $\beta^{2}\propto 1/n$. Since $\|\cdot\|_{1}\leq\sqrt{K}\|\cdot\|_{2}$ (the usual inequality relating $\ell_{1}$ norm to the $\ell_{2}$ norm in $\mathbb{R}^{K}$), this immediately implies that
$\frac{1}{n}\sum_{i\in[n]}\mathbb{E}\left[\mathsf{d}_{\mathsf{TV}}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\right]\leq\frac{K}{4}\cdot\frac{1}{n}\sum_{i\in[n]}\mathbb{E}[\|\hat{w}-\hat{w}^{\setminus i}\|^{2}_{2}]\lesssim K/n,$
or in other words, mean-square stability with respect to $\mathsf{d}_{\mathsf{TV}}$ holds, for $\beta^{2}\propto K/n$. But if the dimension $K$ is large relative to $n$, then this is not a very useful guarantee.
[…]
Problem Setting 2(b): the continuous case
Next, consider a continuous setting, where $\mathcal{A}$ returns a density on $[0,1]$, given by some function $\hat{w}:[0,1]\rightarrow[0,\infty)$ satisfying $\int_{t=0}^{1}\hat{w}(t)\;\mathsf{d}t=1$. In this case we have $\mathsf{d}_{\mathsf{TV}}(w,w^{\prime})=\frac{1}{2}\int_{t=0}^{1}|w(t)-w^{\prime}(t)|\;\mathsf{d}t$.
To formalize this setting, consider the Banach space
$\mathcal{B}=L_{1}([0,1]):=\left\{f:[0,1]\to\mathbb{R}:\int_{t=0}^{1}|f(t)|\;\mathsf{d}t<\infty\right\},$
equipped with the norm $\|f\|=\frac{1}{2}\int_{t=0}^{1}|f(t)|\;\mathsf{d}t$ (note that this differs from the usual $\ell_{1}$ norm by the factor $\frac{1}{2}$, in order to agree with total variation distance). Suppose that our algorithm returns densities that are further constrained to be $L$-Lipschitz, so that we have
$\mathcal{W}=\left\{w:[0,1]\to[0,\infty):\int_{t=0}^{1}w(t)=1,\ \sup_{t\neq t^{\prime}}\frac{|w(t)-w(t^{\prime})|}{|t-t^{\prime}|}\leq L\right\}.$
As for Problem Setting 2(a), we again begin by asking whether the Hilbert space result of Theorem 6 might already be sufficient in this setting. Indeed, while $\mathcal{B}=L_{1}([0,1])$ is not a Hilbert space, the Lipschitz constraint on $\mathcal{W}$ means that $\int_{t=0}^{1}|w(t)|^{2}\;\mathsf{d}t<\infty$ for all $w\in\mathcal{W}$, meaning that $\mathcal{W}$ can be viewed as a subset of the Hilbert space $L_{2}([0,1])$, equipped with norm $\|f\|_{\mathcal{H}}=\|f\|_{L_{2}}=(\int_{t=0}^{1}|f(t)|^{2}\;\mathsf{d}t)^{1/2}$. A straightforward calculation verifies that $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})\propto L^{1/4}$, meaning that by Theorem 6, derandomized bagging satisfies mean-square stability with respect to $\|\cdot\|_{L_{2}}$ for $\beta^{2}\propto L^{1/2}/n$. Since for this function space we have $\|\cdot\|_{L_{1}}\leq\|\cdot\|_{L_{2}}$, this immediately implies that mean-square stability with respect to $\mathsf{d}_{\mathsf{TV}}$ also holds with parameter $\beta^{2}\propto L^{1/2}/n$. However, if the Lipschitz constant $L$ is large, this may not be a satisfying result.
Now we turn to our new result, Theorem 10, to see if the stability guarantee can be improved. First, observe that 2 holds for any $K\geq 1$ if we take $R=1$ and $\rho\propto L/K$. This holds because we can take $v_{j}$ to be the function $v_{j}(t)=K\cdot\mathbf{1}\{\frac{j-1}{K}\leq t\leq\frac{j}{K}\}$, for each $j\in[K]$—then for any $L$-Lipschitz densities $w,w^{\prime}$, the $2L$-Lipschitz difference $w-w^{\prime}$ can be approximated with a piecewise constant function, i.e., a linear combination of these $v_{j}$’s. Choosing $K\propto Ln$, we apply Theorem 10, which then leads to a guarantee on the mean-square stability with respect to $\mathsf{d}_{\mathsf{TV}}$, with $\beta^{2}\propto\frac{\log(Ln)}{n}$. Again, as for the discrete case, if $L$ is large (i.e., $L\gg\log^{2}n$) then this is a much stronger bound than the one we can obtain via the Hilbert space result.
Summary of examples
The following table summarizes our results for the two settings, in terms of the stability guarantee with respect to $\mathsf{d}_{\mathsf{TV}}$. The improved scaling in the right column highlights the utility of Theorem 10 in extending our results to a broader range of settings.
Output space $\mathcal{W}$
Stability guarantee via
Theorem 6 (Hilbert space)
Stability guarantee via
Theorem 10 (Banach space)
Problem Setting 2(a):
distributions on $\{1,\dots,K\}$
$\beta^{2}\propto\frac{K}{n}$ $\beta^{2}\propto\frac{\log K}{n}$
Problem Setting 2(b):
$L$-Lipschitz densities on $[0,1]$
$\beta^{2}\propto\frac{L^{1/2}}{n}$ $\beta^{2}\propto\frac{\log(Ln)}{n}$
Remark 11 (On the necessity of 2).
Our stability guarantee for general Banach spaces essentially requires that $\mathcal{W}$ lies (approximately) in a convex hull of $2K$ many points—namely, $Rv_{1},-Rv_{1}$, …, $Rv_{K},-Rv_{K}$. In particular, this implies that $\mathcal{W}$ is (approximately) contained in a finite-dimensional subspace of $\mathcal{B}$. In contrast, our Hilbert space results apply regardless of dimensionality—while we do require $\mathcal{W}$ to be bounded via its radius, $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})$, there is no implicit cost to the dimension of $\mathcal{H}$ in the results of Theorem 6. It turns out that some assumption of this form is actually required for the Banach case—see 13 (in Section B.2) for an explicit counterexample verifying stability fails without this type of assumption.
6 Experiments
In this section, we study the stability of subbagging in experiments across several simulation and real data settings. Code to reproduce all experiments is available at
6.1 Data and Methods
We start by presenting an overview of our four experiments. Experiments 1, 2, and 3 each examine stability in a Hilbert space, with respect to the appropriate Hilbert norm $\|\cdot\|_{\mathcal{H}}$ (as studied in Section 4). In contrast, the setting of Experiment 4 is a Banach space (as in Section 5).
Experiment 1 (Estimating a regression function using bagged decision trees)
We draw data from the following data-generating process:
[…]
with $n=500$ and $d=40$. We then set $Y_{i}=\frac{\tilde{Y}_{i}-\min_{j}\tilde{Y}_{j}}{\max_{j}\tilde{Y}_{j}-\min_{j}\tilde{Y}_{j}}$ so that each response $Y_{i}$ is in the unit interval $[0,1]$. Note that the algorithm has access to the observed data $\mathcal{D}=(X_{i},Y_{i})_{i\in[n]}$, i.e., $\alpha_{i}$ and $\gamma_{i}$ are latent variables used only to generate the data $\mathcal{D}$. We apply sklearn.tree.DecisionTreeRegressor to train the regression trees, setting max_depth=50 and leaving all other parameters at their default values. We evaluate the stability of the learned function in the $L_{2}$ norm $\|\hat{w}\|_{L_{2}}=\int_{[0,1]^{d}}\hat{w}(x)\textnormal{d}x$.
Experiment 2 (Synthetic controls real data example)
For this experiment, we reproduce Abadie et al. (2015)’s SCM analysis on the 1990 reunification of Germany, as discussed above in Section 1.1 (see Problem Setting 3). The data contains $n=16$ countries (aside from Germany), with $6$ predictors measured for each country (pre-treatment averages of various macroeconomic factors such as inflation and trade openness—see Table 1 of Abadie et al. (2015) for more details). SCM then estimates a weighted combination of these $n$ countries that represents a ‘synthetic West Germany’, meaning that the output of the algorithm is a vector $\hat{w}\in\Delta_{n-1}$. When we run SCM on a subset of the $n=16$ countries, we assign a weight of 00 to each omitted country. We run SCM using the Python package pysyncon and make extensive use of their reproduction of the Abadie et al. (2015) study. We evaluate the stability of these weights in the Euclidean norm $\|\cdot\|_{2}$.
Experiment 3 (Least-squares spectrum analysis)
We study function estimation in a Sobolev space, corresponding to the Hilbert space $W^{s,2}(\mathbb{R})$ with smoothness parameter $s=2$, with the corresponding norm
$\|f\|_{s,2}:=\sqrt{\sum_{k=-\infty}^{\infty}(1+k^{2})^{s}\cdot|\hat{f}(k)|^{2}},$ (3)
where $\hat{f}(k)$ denotes the $k$th Fourier coefficient, indexed by $k\in\mathbb{Z}$. We wish to recover $f$, which is equivalent to estimating the Fourier coefficients. We restrict our estimation problem to even functions, so that we may constrain the Fourier coefficients to be real valued.
Least squares spectral analysis (LSSA) is one method for estimating the Fourier coefficients of a function from irregularly spaced samples Vaníček (1969). Given a data set $\mathcal{D}=(X_{i},Y_{i})_{i\in[n]}$, we first construct a countable set of features $\Psi_{ik}=\exp\left(jkX_{i}\right)$, where $j=\sqrt{-1}$, $i\in[n]$ and $k\in\mathbb{Z}$. The base algorithm $\mathcal{A}(\mathcal{D})$ returns a function $\hat{w}:[-1,1]\to\mathbb{R}$, given by
$\hat{w}(x)=\sum_{k=-\infty}^{\infty}\hat{a}_{k}\exp(jkx),\qquad\textnormal{with}\qquad\hat{a}=\mathop{\rm argmin}_{a\in\mathcal{C}_{n}(R)}\sum_{i=1}^{n}\left(Y_{i}-\sum_{k=-\infty}^{\infty}\Psi_{ik}a_{k}\right)^{2},$
and where the constraint set is given by
$\mathcal{C}_{n}(R):=\left\{(\ldots,a_{-1},a_{0},a_{1},\ldots):a_{k}\in\mathbb{R},a_{k}=0\,\forall\,|k|>\lceil{n/2}\rceil,\sum_{k=-\infty}^{\infty}(1+k^{2})^{2}a_{k}^{2}\leq R^{2}\right\}.$
[…]
Experiment 4 (Stability of the softmax function)
We simulate a binary $n\times d$ data matrix $X\in\{0,1\}^{n\times d}$ using $X_{ij}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\textnormal{Bern}(0.2)$, where $n=2000$ and $d=100$. The base algorithm takes the softmax of the column sums $\mathcal{A}(\mathcal{D}):=\sigma(\mathbf{1}_{n}^{\top}X)\in\Delta_{d-1}$, i.e.
$\mathcal{A}(\mathcal{D})_{j}=\frac{\exp\left(\sum_{i=1}^{n}X_{ij}\right)}{\sum_{k=1}^{d}\exp\left(\sum_{i=1}^{n}X_{ik}\right)}.$
We evaluate stability with respect to the total variation distance $\mathsf{d}_{\mathsf{TV}}$. (As discussed in Section 5, the $\mathsf{d}_{\mathsf{TV}}$ norm is not a Hilbert norm—that is, unlike the first three experiments, stability guarantees for Experiment 4 hold only due to the extension to the Banach space setting given in Theorem 10.)
Experiment 1: Regression trees Refer to caption Experiment 2: Synthetic controls, German reunification example (Abadie et al., 2015) Refer to caption Experiment 3: Least-squares spectral analysis Refer to caption Experiment 4: Softmax Refer to caption
Figure 1: Simulation results comparing the stability of subbagging to that of the corresponding base algorithm under the given norm. Left: Histogram of leave-one-out perturbations. Right: for each $\varepsilon$, the smallest $\delta$ such that the algorithm has tail stability $(\varepsilon,\delta)$ in the sense of Definition 2. Higher curves thus represent greater instability. In the right panels we also plot the stability guarantee of Theorem 6 for Experiments 1-3 and Theorem 10 for Experiment 4 (the guarantees for derandomized bagging). For all four experiments, $m=n/2$ and $B=10,000$.
6.2 Results
Our experiments compare the stability of the base algorithm $\mathcal{A}$ to the stability of subbagging with $m=n/2$ and $B=10,000$. Our results are shown in Figure 1 and Table 2. The left panels of Figure 1 show the histogram of leave-one-out perturbations $\left\|\hat{w}-\hat{w}^{\setminus i}\right\|$ for $i\in\left\{1,\ldots,n\right\}$. In the right panels of Figure 1, for a fixed data set and algorithm, we measure stability by plotting, for each value of $\varepsilon$, the smallest value of $\delta$ such that the algorithm is $\left(\varepsilon,\delta\right)$-stable:
$\delta=\frac{1}{n}\sum_{i=1}^{n}1_{\|\hat{w}-\hat{w}^{\setminus i}\|>\varepsilon}.$
Furthermore, in Table 2, we record the mean-square stability
$\beta^{2}:=\frac{1}{n}\sum_{i=1}^{n}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}.$
For all four experiments, we observe that the subbagged algorithm has better stability than the base algorithm, both in terms of the mean-square stability parameter $\beta^{2}$ and in terms of the tail stability parameters (across all, or nearly all, values of $\varepsilon$).
Setting
Mean-square stability $\beta^{2}$
for base algorithm
Mean-square stability $\beta^{2}$
for subbagged algorithm
Regression trees $0.0107$ $0.0000047$
Synthetic controls (SCM) $0.0532$ $0.0125$
Spectral analysis (LSSA) $40.033$ $0.3017$
Softmax $0.0744$ $0.0014$
Table 2: Results for mean-square stability.
7 Conclusion
In supervised learning, algorithmic stability plays a fundamental role in generalization theory (Bousquet and Elisseeff, 2002; Mukherjee et al., 2006; Shalev-Shwartz et al., 2010). More broadly, various forms of stability are fundamental across applied statistics, in part for their role in establishing reproducibility and interpretability (Yu and Kumbier, 2020; Yu and Barter, 2024). This work provides general, finite sample theoretical guarantees supporting the use of bagging and related techniques as general-purpose stabilizers. Our stability guarantees represent a major generalization of Soloff et al. (2023)—which studied algorithms with outputs on the real line—as our results apply to infinite dimensional settings.
An important remaining open question is how to define and guarantee stability for algorithms with discrete outputs. In the synthetic controls setting, for example, our theory shows that bagging can stabilize the SCM weights. However, our results do not imply stability in the resulting support—that is, the set of control units that receive nonzero support. In a sense, the goal of sparsity is incompatible with stability (Xu et al., 2011), so we need a new framework to reconcile these two desiderata. We leave these questions for future work.
Acknowledgements
We gratefully acknowledge the support of the National Science Foundation via grant DMS-2023109 DMS-2023109. R.F.B. and J.A.S. gratefully acknowledge the support of the Office of Naval Research via grant N00014-20-1-2337. R.M.W. gratefully acknowledges the support of NSF DMS-2235451, Simons Foundation MP-TMPS-00005320, and AFOSR FA9550-18-1-0166.
Appendix A Proofs for the Hilbert space setting
Proof of Theorem 6.
For each $r\in\textnormal{seq}_{[n]}$, define $\hat{w}^{r}=\mathbb{E}_{\xi}[\mathcal{A}(\mathcal{D}^{r};\xi)]$, the expected output of the randomized algorithm for the data set given by $\mathcal{D}^{r}$. From this point on, we write $\mathbb{E}[\cdot]$ to denote expectation with respect to $r\sim\mathcal{Q}_{n}$. We then have $\mathbb{E}[\hat{w}^{r}]=\hat{w}$ and
$\hat{w}^{\setminus i}=\mathbb{E}\left[\hat{w}^{r}\mid i\not\in r\right]=\frac{1}{1-p_{\mathcal{Q}_{n}}}\mathbb{E}\left[\hat{w}^{r}\cdot\mathbf{1}\{i\not\in r\}\right]$
by construction, and therefore, a straightforward calculation shows that
$\hat{w}-\hat{w}^{\setminus i}=\frac{1}{1-p_{\mathcal{Q}_{n}}}\mathbb{E}\left[(\hat{w}^{r}-\hat{w})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right].$ (4)
Therefore,
$\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}_{\mathcal{H}}$ $=\sum_{i\in[n]}\left\langle\hat{w}-\hat{w}^{\setminus i},\frac{1}{1-p_{\mathcal{Q}_{n}}}\mathbb{E}\left[(\hat{w}^{r}-\hat{w})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right]\right\rangle$
$=\frac{1}{1-p_{\mathcal{Q}_{n}}}\mathbb{E}\left[\left\langle\sum_{i\in[n]}(\hat{w}-\hat{w}^{\setminus i})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right),\hat{w}^{r}-\hat{w}\right\rangle\right]$
$\leq\frac{1}{1-p_{\mathcal{Q}_{n}}}\mathbb{E}\left[\left\|\sum_{i\in[n]}(\hat{w}-\hat{w}^{\setminus i})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right\|_{\mathcal{H}}^{2}\right]^{1/2}\cdot\mathbb{E}\left[\left\|\hat{w}^{r}-\hat{w}\right\|_{\mathcal{H}}^{2}\right]^{1/2}$
$\leq\frac{\mathsf{rad}_{\mathcal{H}}(\mathcal{W})}{1-p_{\mathcal{Q}_{n}}}\mathbb{E}\left[\left\|\sum_{i\in[n]}(\hat{w}-\hat{w}^{\setminus i})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right\|_{\mathcal{H}}^{2}\right]^{1/2},$
with the last step holding because we have
$\mathbb{E}\left[\|\hat{w}^{r}-\hat{w}\|_{\mathcal{H}}^{2}\right]=\inf_{w\in\mathcal{W}}\mathbb{E}\left[\|\hat{w}^{r}-w\|_{\mathcal{H}}^{2}\right]\leq\inf_{w\in\mathcal{W}}\sup_{w^{\prime}\in\mathcal{W}}\|w^{\prime}-w\|_{\mathcal{H}}^{2}=\mathsf{rad}^{2}_{\mathcal{H}}(\mathcal{W}),$
where the first equality holds since $\hat{w}=\mathbb{E}[\hat{w}^{r}]$.
To calculate the last remaining expected value, we first note that since we have assumed symmetry and nonpositive covariance for $\mathcal{Q}_{n}$ (see 1), we have $\textnormal{Cov}\left(\mathbf{1}\{i\in r\},\mathbf{1}\{j\in r\}\right)=-q$ for all $i\neq j$, for some $q\geq 0$. Moreover, since the covariance matrix of the random vector $(\mathbf{1}\{i\in r\})_{i\in[n]}$ must be positive semidefinite, we must have $q\leq\frac{p_{\mathcal{Q}_{n}}(1-p_{\mathcal{Q}_{n}})}{n-1}$. We then have
$\mathbb{E}\left[\left\|\sum_{i\in[n]}(\hat{w}-\hat{w}^{\setminus i})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right\|_{\mathcal{H}}^{2}\right]$
$=\sum_{i\in[n]}\textnormal{Var}(\mathbf{1}\{i\in r\})\cdot\|\hat{w}-\hat{w}^{\setminus i}\|^{2}_{\mathcal{H}}+\sum_{i\neq j\in[n]}\textnormal{Cov}\left(\mathbf{1}\{i\in r\},\mathbf{1}\{j\in r\}\right)\cdot\langle\hat{w}-\hat{w}^{\setminus i},\hat{w}-\hat{w}^{\setminus j}\rangle$
$=\sum_{i\in[n]}p_{\mathcal{Q}_{n}}(1-p_{\mathcal{Q}_{n}})\cdot\|\hat{w}-\hat{w}^{\setminus i}\|^{2}_{\mathcal{H}}+\sum_{i\neq j\in[n]}(-q)\cdot\langle\hat{w}-\hat{w}^{\setminus i},\hat{w}-\hat{w}^{\setminus j}\rangle$
$=\left(p_{\mathcal{Q}_{n}}(1-p_{\mathcal{Q}_{n}})+q\right)\cdot\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}_{\mathcal{H}}-q\cdot\left\|\sum_{i\in[n]}\hat{w}-\hat{w}^{\setminus i}\right\|^{2}_{\mathcal{H}}$
$\leq\frac{n}{n-1}\cdot p_{\mathcal{Q}_{n}}(1-p_{\mathcal{Q}_{n}})\cdot\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}_{\mathcal{H}}.$
Combining everything, then, we have
$\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}_{\mathcal{H}}\leq\frac{\mathsf{rad}_{\mathcal{H}}(\mathcal{W})}{1-p_{\mathcal{Q}_{n}}}\left(\frac{n}{n-1}\cdot p_{\mathcal{Q}_{n}}(1-p_{\mathcal{Q}_{n}})\cdot\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}_{\mathcal{H}}\right)^{1/2},$
which yields
$\frac{1}{n}\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}_{\mathcal{H}}\leq\frac{\mathsf{rad}^{2}_{\mathcal{H}}(\mathcal{W})}{n-1}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}$
after rearranging terms. ∎
Proof of Corollary 9.
Let $\hat{w}$ and $\hat{w}^{\setminus i}$ be the output of the bagged algorithm (as in Definition 4), and let $\tilde{w}$ and $\tilde{w}^{\setminus i}$ be the corresponding derandomized bagged estimators (as in Definition 5). For each $i\in[n]$, we have
$\mathbb{E}\left[\|\hat{w}-\hat{w}^{\setminus i}\|^{2}_{\mathcal{H}}\right]=\mathbb{E}\left[\|(\hat{w}-\tilde{w})-(\hat{w}^{\setminus i}-\tilde{w}^{\setminus i})+(\tilde{w}-\tilde{w}^{\setminus i})\|^{2}_{\mathcal{H}}\right]\\ =\mathbb{E}\left[\|(\hat{w}-\tilde{w})-(\hat{w}^{\setminus i}-\tilde{w}^{\setminus i})\|^{2}_{\mathcal{H}}\right]+\|\tilde{w}-\tilde{w}^{\setminus i}\|^{2}_{\mathcal{H}},$
since $\tilde{w}-\tilde{w}^{\setminus i}$ is nonrandom, while $(\hat{w}-\tilde{w})-(\hat{w}^{\setminus i}-\tilde{w}^{\setminus i})$ has mean zero. Next,
$\mathbb{E}\left[\|(\hat{w}-\tilde{w})-(\hat{w}^{\setminus i}-\tilde{w}^{\setminus i})\|^{2}_{\mathcal{H}}\right]\leq 2\mathbb{E}\left[\|\hat{w}-\tilde{w}\|^{2}_{\mathcal{H}}\right]+2\mathbb{E}\left[\|\hat{w}^{\setminus i}-\tilde{w}^{\setminus i}\|^{2}_{\mathcal{H}}\right].$
We next have $\hat{w}=\frac{1}{B}\sum_{b=1}^{B}\mathcal{A}(\mathcal{D}^{r_{b}};\xi_{b})$. Note that the random variables $\mathcal{A}(\mathcal{D}^{r_{b}};\xi_{b})\in\mathcal{H}$ are i.i.d., with mean $\tilde{w}$, and with $\|\mathcal{A}(\mathcal{D}^{r_{b}};\xi_{b})-\tilde{w}\|_{\mathcal{H}}\leq 2\mathsf{rad}_{\mathcal{H}}(\mathcal{W})$ almost surely. Applying Theorem 8,
$\mathbb{E}\left[\|\hat{w}-\tilde{w}\|^{2}_{\mathcal{H}}\right]\leq\frac{4e^{2}\mathsf{rad}^{2}_{\mathcal{H}}(\mathcal{W})}{B}.$
By an identical argument, the same holds for the leave-one-out outputs, $\hat{w}^{\setminus i}$ and $\tilde{w}^{\setminus i}$. Therefore,
$\mathbb{E}\left[\|\hat{w}-\hat{w}^{\setminus i}\|^{2}_{\mathcal{H}}\right]\leq\|\tilde{w}-\tilde{w}^{\setminus i}\|^{2}_{\mathcal{H}}+\frac{16e^{2}\mathsf{rad}^{2}_{\mathcal{H}}(\mathcal{W})}{B}.$
Finally, $\|\tilde{w}-\tilde{w}^{\setminus i}\|^{2}_{\mathcal{H}}$ is bounded by Theorem 6. ∎
Appendix B Proofs and additional results for the Banach space setting
B.1 Proof of main theorem
Proof of Theorem 10.
[…]
$\hat{f}_{i}(\hat{w}-\hat{w}^{\setminus i})=\|\hat{w}-\hat{w}^{\setminus i}\|.$
We then have
$\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}=\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\hat{f}_{i}(\hat{w}-\hat{w}^{\setminus i})$
[…]
$=\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\frac{1}{1-p_{\mathcal{Q}_{n}}}\mathbb{E}\left[\hat{f}_{i}(\hat{w}^{r}-\hat{w})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right]\textnormal{ since $\hat{f}_{i}$ is linear}$
$=\frac{1}{1-p_{\mathcal{Q}_{n}}}\cdot\mathbb{E}\left[\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\hat{f}_{i}(\hat{w}^{r}-\hat{w})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right].$
Next, by 2, for each $r\in\textnormal{seq}_{[n]}$ we can find some $t^{r}=(t^{r}_{1},\dots,t^{r}_{K})\in\mathbb{R}^{K}$ with $\|t^{r}\|_{1}\leq R$ and $\Big{\|}(\hat{w}^{r}-\hat{w})-\sum_{j=1}^{K}t^{r}_{j}v_{j}\Big{\|}\leq\rho$. Then, since $\hat{f}_{i}$ is linear and has unit dual norm $\|\hat{f}_{i}\|_{*}=1$, $\left|\hat{f}_{i}(\hat{w}^{r}-\hat{w})-\sum_{j=1}^{K}t^{r}_{j}\hat{f}_{i}(v_{j})\right|\leq\rho$ for each $i\in[n]$, and so
$\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}$ $=\frac{1}{1-p_{\mathcal{Q}_{n}}}\cdot\mathbb{E}\left[\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\hat{f}_{i}(\hat{w}^{r}-\hat{w})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right]$
$\leq\frac{1}{1-p_{\mathcal{Q}_{n}}}\cdot\mathbb{E}\left[\sum_{j\in[K]}t^{r}_{j}\cdot\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\hat{f}_{i}(v_{j})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right]$
$\hskip 72.26999pt+\frac{\rho}{1-p_{\mathcal{Q}_{n}}}\cdot\mathbb{E}\left[\sum_{i\in[n]}\left|\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right|\right]$
$\leq\frac{R}{1-p_{\mathcal{Q}_{n}}}\cdot\mathbb{E}\left[\max_{j\in[K]}\left|\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\hat{f}_{i}(v_{j})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right|\right]$
$\hskip 72.26999pt+\frac{\rho}{1-p_{\mathcal{Q}_{n}}}\cdot\mathbb{E}\left[\sum_{i\in[n]}\left|\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right|\right],$
where for the last step, we use the fact that $\|t^{r}\|_{1}\leq R$ for any $r$. Since $\mathbb{E}[|\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}|]=2p_{\mathcal{Q}_{n}}(1-p_{\mathcal{Q}_{n}})$ for each $i$, and $\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\leq\sqrt{n}\cdot\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}$, we can rewrite this as
$\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq\frac{R}{1-p_{\mathcal{Q}_{n}}}\cdot$ $\mathbb{E}\left[\max_{j\in[K]}\left|\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\hat{f}_{i}(v_{j})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right|\right]$
$+2\rho\sqrt{n}\cdot p_{\mathcal{Q}_{n}}\cdot\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}.$
Next, define $\hat{p}^{r}=\frac{1}{n}\sum_{i\in[n]}\mathbf{1}\{i\in r\}$, so that $\mathbb{E}[\hat{p}^{r}]=p_{\mathcal{Q}_{n}}$. Using 1 we can verify that $\textnormal{Var}(\hat{p}^{r})\leq\frac{p_{\mathcal{Q}_{n}}(1-p_{\mathcal{Q}_{n}})}{n}$. Then, since $|\hat{f}_{i}(v_{j})|\leq 1$ for all $i,j$,
$\mathbb{E}\left[\max_{j\in[K]}\left|\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\hat{f}_{i}(v_{j})\cdot\left(\hat{p}^{r}-p_{\mathcal{Q}_{n}}\right)\right|\right]=\mathbb{E}\left[\max_{j\in[K]}\left|\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\hat{f}_{i}(v_{j})\right|\cdot|\hat{p}^{r}-p_{\mathcal{Q}_{n}}|\right]\\ \leq\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\mathbb{E}\left[|\hat{p}^{r}-p_{\mathcal{Q}_{n}}|\right]\leq\sqrt{n}\cdot\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}\cdot\sqrt{\textnormal{Var}(\hat{p}^{r})}\\ \leq\sqrt{p_{\mathcal{Q}_{n}}(1-p_{\mathcal{Q}_{n}})}\cdot\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}.$
Combining with the work above, then,
$\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq\frac{R}{1-p_{\mathcal{Q}_{n}}}\cdot\mathbb{E}\left[\max_{j\in[K]}\left|\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\hat{f}_{i}(v_{j})\cdot\left(\mathbf{1}\{i\in r\}-\hat{p}^{r}\right)\right|\right]\\ {}+R\sqrt{\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}\cdot\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}+2\rho\sqrt{n}\cdot p_{\mathcal{Q}_{n}}\cdot\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}.$
Now we bound the remaining expected value. Let $c^{(j)}\in\mathbb{R}^{n}$ have entries $c^{(j)}_{i}=\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\hat{f}_{i}(v_{j})$, and observe that $\|c^{(j)}\|^{2}_{2}\leq\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}$ and $\|c^{(j)}\|_{\infty}\leq\max_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|$, for each $j$, because $|\hat{f}_{i}(v_{j})|\leq 1$. Moreover, for any $i$, $\hat{w}-\hat{w}^{\setminus i}=p_{\mathcal{Q}_{n}}(\hat{w}^{i}-\hat{w}^{\setminus i})$, where $\hat{w}^{i}=\mathbb{E}[\hat{w}^{r}\mid i\in r]$, and so $\|\hat{w}-\hat{w}^{\setminus i}\|\leq p_{\mathcal{Q}_{r}}R$ (by 2), and so $\|c^{(j)}\|_{\infty}\leq p_{\mathcal{Q}_{n}}R$ for all $j$. Applying Lemma 12 (with vectors $c^{(1)},\dots,c^{(K)},-c^{(1)},\dots,-c^{(K)}$), then,
$\mathbb{E}\left[\max_{j\in[K]}\left|\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\hat{f}_{i}(v_{j})\cdot\left(\mathbf{1}\{i\in r\}-\hat{p}^{r}\right)\right|\,\middle|\,\hat{p}^{r}\right]\\ \leq\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}\cdot\sqrt{2(1+\eta_{n})(\hat{p}^{r}(1-\hat{p}^{r})+4\eta_{n})\log(2K)}+\frac{2}{3}(1+\eta_{n})\log(2K)\cdot p_{\mathcal{Q}_{n}}R.$
Finally, since $t\mapsto\sqrt{t(1-t)+c}$ is concave for $t\in[0,1]$ and for fixed $c\geq 0$, marginalizing over $\hat{p}^{r}$, and applying Jensen’s inequality, yields
$\mathbb{E}$ $\left[\max_{j\in[K]}\left|\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|\cdot\hat{f}_{i}(v_{j})\cdot\left(\mathbf{1}\{i\in r\}-\hat{p}^{r}\right)\right|\right]$
$\leq\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}\cdot\sqrt{2(1+\eta_{n})(p_{\mathcal{Q}_{n}}(1-p_{\mathcal{Q}_{n}})+4\eta_{n})\log(2K)}+\frac{2}{3}(1+\eta_{n})\log(2K)\cdot p_{\mathcal{Q}_{n}}R.$
Combining all the above calculations proves that $\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}\cdot A+B$, where
$A=R\left[\sqrt{2(1+\eta_{n})\left(\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{4\eta_{n}}{(1-p_{\mathcal{Q}_{n}})^{2}}\right)\log(2K)}+\sqrt{\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}\right]+2\rho\sqrt{n}\cdot p_{\mathcal{Q}_{n}}$
and
$B=R^{2}\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}\cdot\frac{2}{3}(1+\eta_{n})\log(2K)\leq\frac{A^{2}}{3}.$
Then, solving the quadratic,
$\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}\cdot A+\frac{A^{2}}{3}\ \Longrightarrow\ \sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq A^{2}\cdot\left(\frac{1+\sqrt{7/3}}{2}\right)^{2}.$
This completes the proof of the theorem. ∎
Lemma 12.
Let $B\in\{0,1\}^{n}$ be a uniformly random permutation of $(\mathbf{1}_{k},\mathbf{0}_{n-k})$, and let $p=k/n$. Let $c^{(1)},\dots,c^{(K)}\in\mathbb{R}^{n}$ be fixed, with $\max_{j\in[K]}\|c^{(j)}\|_{2}\leq C_{2}$ and $\max_{j\in[K]}\|c^{(j)}\|_{\infty}\leq C_{\infty}$. Then
$\mathbb{E}\left[\max_{j\in[K]}c^{(j)\top}(B-p\mathbf{1}_{n})\right]\leq C_{2}\sqrt{2(1+\eta_{n})(p(1-p)+4\eta_{n})\log K}+\frac{2C_{\infty}}{3}(1+\eta_{n})\log K,$
where $\eta_{n}=\frac{H_{n}-1}{n-H_{n}}$, for $H_{n}=1+\frac{1}{2}+\dots+\frac{1}{n}$ denoting the $n$th harmonic number.
Proof.
Applying (Barber, 2024, Theorem 10), for each $j$, for any $0<\lambda<\frac{3}{2\|c^{(j)}\|_{\infty}(1+\eta_{n})}$, it holds that
$\mathbb{E}\left[\exp\left\{\lambda c^{(j)\top}(B-p\mathbf{1}_{n})\right\}\right]\leq\exp\left\{\frac{\lambda^{2}\|c^{(j)}\|^{2}_{2}(1+\eta_{n})}{2(1-\frac{2\lambda}{3}\|c^{(j)}\|_{\infty}(1+\eta_{n}))}\cdot\left(p(1-p)+4\eta_{n}\right)\right\}.$
Therefore, from the definition of $C_{2}$ and $C_{\infty}$, for any $0<\lambda<\frac{3}{2(1+\eta_{n})C_{\infty}}$, this implies
$\mathbb{E}\left[\exp\left\{\lambda c^{(j)\top}(B-p\mathbf{1}_{n})\right\}\right]\leq\exp\left\{\frac{\lambda^{2}C_{2}^{2}(1+\eta_{n})}{2(1-\frac{2\lambda}{3}(1+\eta_{n})C_{\infty})}\cdot\left(p(1-p)+4\eta_{n}\right)\right\}$
for each $j$. We then have
[…]
$\leq\lambda^{-1}\log\left(\sum_{j\in[K]}\mathbb{E}\left[\exp\left\{\lambda c^{(j)\top}(B-p\mathbf{1}_{n})\right\}\right]\right)$
$\leq\lambda^{-1}\log\left(K\cdot\exp\left\{\frac{\lambda^{2}C_{2}^{2}(1+\eta_{n})}{2(1-\frac{2\lambda}{3}(1+\eta_{n})C_{\infty})}\cdot\left(p(1-p)+4\eta_{n}\right)\right\}\right)$
$=\frac{\log K}{\lambda}+\frac{\lambda C_{2}^{2}(1+\eta_{n})}{2(1-\frac{2\lambda}{3}(1+\eta_{n})C_{\infty})}\cdot\left(p(1-p)+4\eta_{n}\right).$
Taking
$\lambda=\frac{\sqrt{\frac{2(1+\eta_{n})\log K}{C_{2}^{2}(p(1-p)+4\eta_{n})}}}{(1+\eta_{n})\left(1+\frac{2C_{\infty}}{3}\sqrt{\frac{2(1+\eta_{n})\log K}{C_{2}^{2}(p(1-p)+4\eta_{n})}}\right)}<\frac{3}{2(1+\eta_{n})C_{\infty}},$
this yields the desired bound. ∎
B.2 Is 2 needed?
We return to the role of 2, as discussed in Remark 11. In particular, this assumption requires that $\mathcal{W}$ is approximately contained in a $K$-dimensional subspace of $\mathcal{B}$.
We will now see that no uniform stability guarantee can hold without this type of condition. A counterexample can be found with a standard infinite-dimensional Banach space: $\mathcal{B}=L_{1}(\mathbb{N})=\{(x_{1},x_{2},\dots):\sum_{j\geq 1}|x_{j}|<\infty\}$. We will equip $\mathcal{B}$ with the norm induced by the $\mathsf{d}_{\mathsf{TV}}$ metric, given by $\|x\|=\frac{1}{2}\|x\|_{1}=\frac{1}{2}\sum_{j\geq 1}|x_{j}|$. This is an infinite-dimensional and separable Banach space.
[…]
Proposition 13.
Let $\mathcal{Z}=\mathbb{N}$, the set of positive integers. Let $\mathcal{B}=L_{1}(\mathbb{N})$ (equipped with the norm $\|x\|=\frac{1}{2}\sum_{i\geq 1}|x_{i}|$), and let $\mathcal{W}=\Delta_{\infty}$. Then there exists a (deterministic) algorithm $\mathcal{A}:\cup_{n\geq 0}\mathcal{Z}^{n}\times[0,1]\to\mathcal{W}$, such that for any distribution $\mathcal{Q}_{n}$ satisfying 1, derandomized bagging fails to satisfy mean-square stability with respect to $\mathsf{d}_{\mathsf{TV}}$ for any $\beta<p$.
In other words, stability results of the type shown in Theorem 10 guarantee a bound on $\beta^{2}$ (the stability parameter), uniformly over all algorithms $\mathcal{A}$, that vanishes as $n\to\infty$. In contrast, the above example, we show that—for a particular construction of the algorithm $\mathcal{A}$—the stability parameter $\beta^{2}$ is bounded away from zero even as $n\to\infty$.
Proof of 13.
Let $\mathcal{S}$ denote all finite subsets of $\mathbb{N}$, and let $\psi:\mathcal{S}\to\mathbb{N}$ be a bijective map that assigns each finite subset of $\mathbb{N}$ to a (unique) positive integer. For any dataset $\mathcal{D}=(Z_{1},\dots,Z_{m})\in\mathbb{N}^{m}$, we define
$S(\mathcal{D})=\left\{i\in\mathbb{N}:\sum_{j\in[m]}\mathbf{1}\{Z_{j}=i\}\geq 1\right\},$
the set of values appearing in the data set. Then define the algorithm as $\mathcal{A}(\mathcal{D};\xi)=\mathbf{e}_{\psi(S(\mathcal{D}))}\in\mathcal{W},$ where $\mathbf{e}_{j}$ is the $j$th canonical basis vector, with a $1$ in the $j$th location and 00’s elsewhere.
Now consider data set $\mathcal{D}=(1,\dots,n)$. Let $\hat{w}$ and $\hat{w}^{\setminus i}$ denote the derandomized bagged estimators, as before, for data sets $\mathcal{D}$ and $\mathcal{D}^{\setminus i}$, respectively. By construction of $\mathcal{A}$, we can verify that $\hat{w}_{j}=\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{\psi(S(\mathcal{D}^{r}))=j\}$ and similarly $\hat{w}^{\setminus i}_{j}=\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{\psi(S(\mathcal{D}^{r}))=j\mid i\not\in r\}$. Therefore,
$\|\hat{w}-\hat{w}^{\setminus i}\|$ $=\frac{1}{2}\sum_{j\in\mathbb{N}}\left|\hat{w}_{j}-\hat{w}^{\setminus i}_{j}\right|=\frac{1}{2}\sum_{j\in\mathbb{N}}\left|\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{\psi(S(\mathcal{D}^{r}))=j\}-\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{\psi(S(\mathcal{D}^{r}))=j\mid i\not\in r\}\right|$
$=\frac{1}{2}\sum_{j\in\mathbb{N}}\left|\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{\psi(S(\mathcal{D}^{r}))=j\}-\frac{\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{\psi(S(\mathcal{D}^{r}))=j\}}{1-p_{\mathcal{Q}_{n}}}\cdot\mathbf{1}\{i\not\in\psi^{-1}(j)\}\right|$
$=\frac{1}{2}\sum_{j\in\mathbb{N}}\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{\psi(S(\mathcal{D}^{r}))=j\}\cdot\left|1-\frac{\mathbf{1}\{i\not\in\psi^{-1}(j)\}}{1-p_{\mathcal{Q}_{n}}}\right|$
$=\frac{1}{2}\sum_{j\in\mathbb{N}}\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{\psi(S(\mathcal{D}^{r}))=j\}\cdot\frac{(1-p_{\mathcal{Q}_{n}})\cdot\mathbf{1}\{i\in S(\mathcal{D}^{r})\}+p_{\mathcal{Q}_{n}}\cdot\mathbf{1}\{i\not\in S(\mathcal{D}^{r})\}}{1-p_{\mathcal{Q}_{n}}}$
$=\frac{1}{2}\frac{1}{1-p_{\mathcal{Q}_{n}}}\left((1-p_{\mathcal{Q}_{n}})\cdot\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{i\in S(\mathcal{D}^{r})\}+p_{\mathcal{Q}_{n}}\cdot\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{i\not\in S(\mathcal{D}^{r})\}\right)$
$=p_{\mathcal{Q}_{n}},$
since $\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{i\in(S(\mathcal{D}^{r}))\}=\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{i\in r\}=p_{\mathcal{Q}_{n}}$. Since this holds for every $i$, we have $\frac{1}{n}\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\backslash i}\|^{2}=p_{\mathcal{Q}_{n}}^{2}$, which completes the proof. ∎
B.2.1 Remark on dimensionality
To relate the construction in this proof back to 2 and Theorem 10, note that in the construction of $\mathcal{A}$, at each sample size $n$ we are essentially working in a $K$-dimensional space for $K=2^{n}$—that is, even though $\mathcal{W}=\Delta_{\infty}$ is the space of all infinite sequences, the algorithm $\mathcal{A}$ only ever returns sequences $\hat{w}$ with nonzero values in the first $K=2^{n}$ entries (and in particular, $\hat{w}$ will always lie in the convex hull of $\mathbf{e}_{1},\dots,\mathbf{e}_{K}$).
Compare this to the result of Theorem 10, which essentially suggests that mean-square stability vanishes at the rate $\beta^{2}\propto\frac{\log K}{n}$ if $\mathcal{W}$ can be approximated as a convex hull of $\mathcal{O}(K)$ many points. Since we have taken $\frac{\log K}{n}$ to be nonvanishing in this example (due to $K=2^{n}$ being the “effective dimension”, as described above), stability is no longer ensured as $n\to\infty$.
B.3 Additional results
First, we state a result for tail stability in the Banach space setting.
Corollary 14.
[…]
$\delta\epsilon^{2}\geq 1.6\left[R\sqrt{\frac{2(1+\eta_{n})\log(2K)}{n}\left(\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{4\eta_{n}}{(1-p_{\mathcal{Q}_{n}})^{2}}\right)}+R\sqrt{\frac{1}{n}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}+2\rho\cdot p_{\mathcal{Q}_{n}}\right]^{2}.$
Proof.
This result follows immediately from Theorem 10 together with 3. ∎
Finally, we state a finite-$B$ result for the Banach space setting.
Corollary 15.
Under the notation and assumptions of Theorem 10, let $B\geq 1$ be fixed. Then bagging, run with base algorithm $\mathcal{A}$ and resampling distribution $\mathcal{Q}_{n}$ and with $B$ bags, has mean-square stability $\beta^{2}$ with respect to $\|\cdot\|_{\mathcal{H}}$, with
$\beta^{2}:=\inf_{\varepsilon>0}\Bigg{\{}(1+\epsilon)\cdot 1.6\bigg{[}R\sqrt{\frac{2(1+\eta_{n})\log(2K)}{n}\left(\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{4\eta_{n}}{(1-p_{\mathcal{Q}_{n}})^{2}}\right)}+R\sqrt{\frac{1}{n}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}\\ {}+2\rho\cdot p_{\mathcal{Q}_{n}}\bigg{]}^{2}{}+(1+\epsilon^{-1})\cdot\left(\frac{8KR^{2}}{B}+8\rho^{2}\right)\Bigg{\}}$
Proof.
Let $\hat{w}$ and $\hat{w}^{\setminus i}$ be the output of the bagged algorithm (as in Definition 4), and let $\tilde{w}$ and $\tilde{w}^{\setminus i}$ be the corresponding derandomized bagged estimators (as in Definition 5). We have
$\frac{1}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\right]=\frac{1}{n}\sum_{i\in[n]}\mathbb{E}\left[\|(\hat{w}-\tilde{w})-(\hat{w}^{\setminus i}-\tilde{w}^{\setminus i})+(\tilde{w}-\tilde{w}^{\setminus i})\|^{2}\right]\\ \leq\frac{2(1+\epsilon^{-1})}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}-\tilde{w}\|^{2}\right]+\frac{2(1+\epsilon^{-1})}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}^{\setminus i}-\tilde{w}^{\setminus i}\|^{2}\right]+\frac{1+\epsilon}{n}\sum_{i\in[n]}\|\tilde{w}-\tilde{w}^{\setminus i}\|^{2},$
since $(a+b+c)^{2}\leq(1+\epsilon)a^{2}+(1+\epsilon^{-1})(b+c)^{2}\leq(1+\epsilon)a^{2}+2(1+\epsilon^{-1})(b^{2}+c^{2})$, for any $a,b,c\geq 0$. The last term in the upper bound is bounded by Theorem 10. Now we bound the remaining terms.
First, let $\hat{w}^{\prime}$ be an i.i.d. copy of $\hat{w}$. Under the same notation as in the proof of Theorem 10, let $\|\cdot\|_{*}$ denote the dual norm. Then
$\mathbb{E}[\|\hat{w}-\tilde{w}\|^{2}]=\mathbb{E}\left[\sup_{f:\|f\|_{*}\leq 1}f(\hat{w}-\tilde{w})^{2}\right]=\mathbb{E}\left[\sup_{f:\|f\|_{*}\leq 1}f\left(\hat{w}-\mathbb{E}[\hat{w}^{\prime}\mid\hat{w}]\right)^{2}\right]\\ =\mathbb{E}\left[\sup_{f:\|f\|_{*}\leq 1}\mathbb{E}\left[f\left(\hat{w}-\hat{w}^{\prime}\right)\mid\hat{w}\right]^{2}\right]\leq\mathbb{E}\left[\mathbb{E}\left[\sup_{f:\|f\|_{*}\leq 1}f\left(\hat{w}-\hat{w}^{\prime}\right)^{2}\mid\hat{w}\right]\right]=\mathbb{E}[\|\hat{w}-\hat{w}^{\prime}\|^{2}],$
where the second step holds since $\hat{w}\perp\!\!\!\perp\hat{w}^{\prime}$ and $\mathbb{E}[\hat{w}^{\prime}]=\mathbb{E}[\hat{w}]=\tilde{w}$. Now let $\hat{w}^{(b)}=\mathcal{A}(\mathcal{D}^{r_{b}};\xi_{b})$, for each $b\in[B]$, so that we have $\hat{w}=\frac{1}{B}\sum_{b=1}^{B}\hat{w}^{(b)}$, and similarly write $\hat{w}^{\prime}{}^{(b)}=\mathcal{A}(\mathcal{D}^{r_{b}^{\prime}};\xi_{b}^{\prime})$, where the $r_{b}^{\prime}$ and $\xi^{\prime}_{b}$ samples are independently drawn, so that $\hat{w}^{\prime}=\frac{1}{B}\sum_{b=1}^{B}\hat{w}^{\prime}{}^{(b)}$. Then $\hat{w}-\hat{w}^{\prime}$ is equal in distribution to $\frac{1}{B}\sum_{b=1}^{B}\zeta_{b}\left(\hat{w}^{(b)}-\hat{w}^{\prime}{}^{(b)}\right)$ where the $\zeta_{b}$’s are i.i.d. random signs.
Now we will bound $\mathbb{E}[\|\frac{1}{B}\sum_{b=1}^{B}\zeta_{b}(w_{b}-w^{\prime}_{b})\|^{2}]$ for any fixed $w_{1},\dots,w_{B},w^{\prime}_{1},\dots,w^{\prime}_{B}\in\mathcal{W}$. For each $b$ we can write $w_{b}-w^{\prime}_{b}=\sum_{j=1}^{K}t^{b}_{j}v_{j}+\Delta_{b}$ with $\|t^{b}\|_{1}\leq R$ and $\|\Delta_{b}\|\leq\rho$, by 2. Then
$\left\|\frac{1}{B}\sum_{b=1}^{B}\zeta_{b}(w_{b}-w^{\prime}_{b})\right\|=\left\|\frac{1}{B}\sum_{b=1}^{B}\zeta_{b}\left(\sum_{j=1}^{K}t^{b}_{j}v_{j}+\Delta_{b}\right)\right\|\\ \leq\sum_{j=1}^{K}\left\|\frac{1}{B}\sum_{b=1}^{B}\zeta_{b}t^{b}_{j}v_{j}\right\|+\left\|\frac{1}{B}\sum_{b=1}^{B}\zeta_{b}\Delta_{b}\right\|\leq\sum_{j=1}^{K}\left|\frac{1}{B}\sum_{b=1}^{B}\zeta_{b}t^{b}_{j}\right|+\rho,$
since $\|v_{j}\|=1$ for all $j$, and $\|\Delta_{b}\|\leq\rho$ for all $b$. Therefore,
$\mathbb{E}\left[\left\|\frac{1}{B}\sum_{b=1}^{B}\zeta_{b}(w_{b}-w^{\prime}_{b})\right\|^{2}\right]\leq\mathbb{E}\left[\left(\sum_{j=1}^{K}\left|\frac{1}{B}\sum_{b=1}^{B}\zeta_{b}t^{b}_{j}\right|+\rho\right)^{2}\right]\leq 2K\sum_{j=1}^{K}\mathbb{E}\left[\left|\frac{1}{B}\sum_{b=1}^{B}\zeta_{b}t^{b}_{j}\right|^{2}\right]+2\rho^{2}\\ \leq\frac{2K}{B^{2}}\sum_{j=1}^{K}\sum_{b=1}^{B}(t_{j}^{b})^{2}+2\rho^{2}=\frac{2K}{B^{2}}\sum_{b=1}^{B}\|t^{b}\|^{2}_{2}+2\rho^{2}\leq\frac{2K}{B^{2}}\sum_{b=1}^{B}\|t^{b}\|^{2}_{1}+2\rho^{2}\leq\frac{2KR^{2}}{B}+2\rho^{2}.$
Combining all our work we then have
$\mathbb{E}[\|\hat{w}-\tilde{w}\|^{2}]\leq\mathbb{E}[\|\hat{w}-\hat{w}^{\prime}\|^{2}]=\mathbb{E}\left[\left\|\frac{1}{B}\sum_{b=1}^{B}\zeta_{b}(\hat{w}^{b}-\hat{w}^{\prime}{}^{b})\right\|^{2}\right]\\ =\mathbb{E}\left[\mathbb{E}\left[\left\|\frac{1}{B}\sum_{b=1}^{B}\zeta_{b}(\hat{w}^{b}-\hat{w}^{\prime}{}^{b})\right\|^{2}\ \middle|\ \hat{w}^{1},\dots,\hat{w}^{B},\hat{w}^{\prime}{}^{1},\dots,\hat{w}^{\prime}{}^{B}\right]\right]\leq\frac{2KR^{2}}{B}+2\rho^{2}.$
The same bound holds for $\mathbb{E}[\|\hat{w}^{\setminus i}-\tilde{w}^{\setminus i}\|^{2}]$. Combining all our calculations establishes the result. ∎

### v2

@@
 In Bayesian data analysis, given a data set $\mathcal{D}$ containing $n$ observations drawn from some distribution $p_{\theta}$, we are interested in the posterior distribution of the parameter $\theta$.
-For instance, $\theta$ might represent a particular model from a finite set of possibilities, in which case the posterior distribution on $\theta$ is simply a vector of weights on the finite set of possibilities.
-Alternatively, $\theta$ might take values in $\mathbb{R}^{d}$ (e.g., coefficients in some regression model), in which case the posterior distribution is a distribution over $\mathbb{R}^{d}$—a more complex object.
+For instance, $\theta$ might represent a particular model from a finite set of possibilities, in which case the posterior distribution on $\theta$ is simply a vector of weights on this finite set.
+Alternatively, $\theta$ might take values in a continuous space such as $[0,1]$ (e.g., if $\theta$ represents a probability) or $\mathbb{R}^{d}$ (e.g., coefficients for some regression model), in which case the posterior distribution is a distribution over $[0,1]$ or over $\mathbb{R}^{d}$—a more complex object.
 Bayesian inference may sometimes lead to answers that are highly sensitive with respect to perturbations of the data.
@@
 $\ell_{q}$-ball $\mathbb{B}_{q}^{d}(R)$ $\ell_{q}$-norm $\|\cdot\|_{q}$
-Density estimation $L$-Lipschitz densities on $[0,1]$ Total variation dist. $\mathsf{d}_{\mathsf{TV}}$
+Density estimation
+(e.g., Bayesian posterior)
+$L$-Lipschitz densities on $[0,1]$ Total variation dist. $\mathsf{d}_{\mathsf{TV}}$
 Nonparametric regression Sobolev ball of functions $\mathbb{R}^{d}\to\mathbb{R}$ Sobolev norm $\|\cdot\|_{s,2}$
-Table 1: Example problem settings for our stability framework. The first three examples correspond to Problem Settings 1-3 described in Section 1.
+Table 1: Example problem settings for our stability framework, including the examples highlighted in Problem Settings 1–3 (in Section 1).
 2 Framework: algorithmic stability in a metric space
@@
 $\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\xi}\left[\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\right]\leq\beta^{2}.$
-where $\hat{w}=\mathcal{A}(\mathcal{D};\xi)$ is the output of the algorithm run on the full data set $\mathcal{D}$, $\hat{w}^{\setminus i}=\mathcal{A}(\mathcal{D}^{\setminus i};\xi)$ is the output of the algorithm run on the leave-one-out data set $\mathcal{D}^{\setminus i}=(Z_{j})_{j\neq i}$ and the expectation averages over $\xi\sim\text{Uniform}[0,1]$.
+where $\hat{w}=\mathcal{A}(\mathcal{D};\xi)$ is the output of the algorithm run on the full data set $\mathcal{D}$, $\hat{w}^{\setminus i}=\mathcal{A}(\mathcal{D}^{\setminus i};\xi)$ is the output of the algorithm run on the leave-one-out data set $\mathcal{D}^{\setminus i}=(Z_{j})_{j\neq i}$, and the expectation averages over $\xi\sim\text{Uniform}[0,1]$.
 This first definition captures the average perturbation in $\hat{w}$ when one data point is removed at random from some data set $\mathcal{D}\in\mathcal{Z}^{n}$—that is, $\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)$, averaged over which data point $i$ gets removed.
@@
 An algorithm $\mathcal{A}$ has tail stability $(\varepsilon,\delta)$ with respect to a metric $\mathsf{d}$ if, for all data sets $\mathcal{D}\in\mathcal{Z}^{n}$ of size $n$,
-$\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}(\hat{w},\hat{w}^{\setminus i})>\varepsilon\right\}\leq\delta,$
+$\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}(\hat{w},\hat{w}^{\setminus i})\geq\varepsilon\right\}\leq\delta,$
 where $\hat{w}=\mathcal{A}(\mathcal{D};\xi)$ and $\hat{w}^{\setminus i}=\mathcal{A}(\mathcal{D}^{\setminus i};\xi)$ as in Definition 1.
@@
 The first claim follows by Markov’s inequality, since
-$\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)>\varepsilon\right\}=\mathbb{E}_{\xi}\left[\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)>\varepsilon^{2}\right\}\right]\\ \leq\varepsilon^{-2}\mathbb{E}_{\xi}\left[\frac{1}{n}\sum_{i=1}^{n}\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\right]\leq\varepsilon^{-2}\beta^{2}.$
-For the second claim, since $\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\leq\varepsilon^{2}+\mathbf{1}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)>\varepsilon\right\}\cdot\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right)$, we have
-$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\xi}\,\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)$ $\leq\varepsilon^{2}+\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)>\varepsilon\right\}\cdot\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right)$
+$\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)\geq\varepsilon\right\}$ $=\mathbb{E}_{\xi}\left[\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\geq\varepsilon^{2}\right\}\right]$
+$\leq\varepsilon^{-2}\mathbb{E}_{\xi}\left[\frac{1}{n}\sum_{i=1}^{n}\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\right]\leq\varepsilon^{-2}\beta^{2}.$
+For the second claim, since $\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\leq\varepsilon^{2}+\mathbf{1}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)\geq\varepsilon\right\}\cdot\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right)$, we have
+$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\xi}\left[\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\right]$ $\leq\varepsilon^{2}+\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)\geq\varepsilon\right\}\cdot\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right)$
 $\leq\varepsilon^{2}+\delta\cdot\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right),$
@@
 We now define some notation that unifies these two variants (as well as allowing for other resampling schemes, as we discuss below).
-Define the set of finite sequences of indices in $[n]$:
+Following the framework and notation of Soloff et al. (2024), we define the set of finite sequences of indices in $[n]$:
 $\textnormal{seq}_{[n]}:=\{(i_{1},\ldots,i_{k}):k\geq 0,\,i_{1},\ldots,i_{k}\in[n]\}.$
@@
 To study the stability of the bagged algorithm (Definition 4), it is not sufficient to consider only $\mathcal{Q}_{n}$, the distribution on $\textnormal{seq}_{[n]}$—when we compare to the result of the algorithm on the leave-one-out data set $\mathcal{D}^{\setminus i}$, we now have sample size $n-1$ (instead of $n$), and consequently we need to specify a distribution $\mathcal{Q}_{n-1}$ on $\textnormal{seq}_{[n-1]}$, as well.
-The four conditions of 2 are all satisfied by the standard resampling schemes described above, with $p_{\mathcal{Q}_{n}}=1-\left(1-\frac{1}{n}\right)^{m}$ for bagging (sampling $m$ out of $n$ points uniformly with replacement), and $p_{\mathcal{Q}_{n}}=\frac{m}{n}$ for subbagging (sampling $m$ out of $n$ points uniformly without replacement).
+The four conditions of 2 are all satisfied by the standard resampling schemes described above, with $p_{\mathcal{Q}_{n}}=1-\left(1-\frac{1}{n}\right)^{m}$ for bootstrapping (sampling $m$ out of $n$ points uniformly with replacement), and $p_{\mathcal{Q}_{n}}=\frac{m}{n}$ for subbagging (sampling $m$ out of $n$ points uniformly without replacement).
 4 Main results for a Hilbert space
@@
   • If $\mathcal{W}\subset\mathbb{R}^{d}$, then the Hilbert norm may be chosen to be the Euclidean norm $\|\cdot\|_{2}$, or more generally, a Mahalanobis distance $\|w\|_{\mathcal{H}}=\sqrt{w^{\top}Sw}$ for a fixed positive definite matrix $S$.
-  • If $\mathcal{W}\subset L_{2}(\mathbb{R}^{d})$, the space of square-integrable functions $\mathbb{R}^{d}\to\mathbb{R}$, then the Hilbert norm is given by the $L_{2}$ norm on functions, $\|w\|_{\mathcal{H}}=(\int_{\mathbb{R}^{d}}w(x)^{2}\;\mathsf{d}x)^{1/2}$.
-Alternatively, we can take a Sobolev space of functions $\mathbb{R}^{d}\to\mathbb{R}$ with smoothness parameter $s$ with the Hilbert norm being the Sobolev norm $\|\cdot\|_{s,2}$ in (3); see Experiment 3 in Section 6.1 for details.
+  • If $\mathcal{W}\subset L_{2}(\mathbb{R}^{d})$, the space of square-integrable functions $\mathbb{R}^{d}\to\mathbb{R}$, then the Hilbert norm is given by the $L_{2}$ norm on functions, $\|w\|_{\mathcal{H}}=(\int_{\mathbb{R}^{d}}w(x)^{2}\;\textnormal{d}x)^{1/2}$.
+Alternatively, we can take a Sobolev space of functions $\mathbb{R}^{d}\to\mathbb{R}$ with smoothness parameter $s$ with the Hilbert norm being the Sobolev norm $\|\cdot\|_{s,2}$ (defined formally later on in (3)—see Experiment 3 in Section 6.1 for details).
 Theorem 6.
@@
 $\delta\varepsilon^{2}\geq\frac{\mathsf{rad}_{\mathcal{H}}^{2}(\mathcal{W})}{n-1}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}.$ (2)
-Corollary 7 generalizes the main result of Soloff et al.
-(2023), which focuses on the supervised learning setting where the data points $Z_{i}=(X_{i},Y_{i})$ contain features $X_{i}\in\mathcal{X}$ and a response $Y_{i}\in\mathbb{R}$, and the output of the algorithm $\mathcal{A}$ is a predicted value $\hat{f}(x)$, at a test point $x\in\mathcal{X}$—i.e., in that work, the output space $\mathcal{H}=\mathbb{R}$.
 Comparing to prior bounds
-The main result of (Soloff et al., 2023, Theorem 8), ensuring the stability of algorithms that return a real-valued output, can be recovered as a special case of Corollary 7 above, by taking $\mathcal{W}=[0,1]\subseteq\mathcal{H}=\mathbb{R}$ (note that we then have $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})=\frac{1}{2}$).
+The main result of (Soloff et al., 2024, Theorem 8), ensuring the stability of algorithms that return a real-valued output, can be recovered as a special case of Corollary 7 above, by taking $\mathcal{W}=[0,1]\subseteq\mathcal{H}=\mathbb{R}$ (note that we then have $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})=\frac{1}{2}$).
 But remarkably, the result of Theorem 6 shows that there is no price to pay for high-dimensionality (or even for infinite dimensionality): the stability guarantee is exactly the same for any Hilbert space, regardless of dimension, as long as the radius of the output space $\mathcal{W}$ is bounded.
@@
 Since we are simply taking a Monte Carlo approximation to the derandomized bagged estimator of Definition 5, we would expect to incur error on the order of $\frac{1}{\sqrt{B}}$.
-To formalize this, we first state a generalization of the Azuma-Hoeffding inequality:
+To formalize this, we first recall a generalization of the Azuma-Hoeffding inequality:
 Theorem 8 ((Hayes, 2005)).
-Let $\hat{w}_{1},\ldots,\hat{w}_{B}\in\mathcal{H}$ be independent random variables in a Hilbert space $\mathcal{H}$ such that $\|\hat{w}_{b}-\mathbb{E}\hat{w}_{b}\|_{\mathcal{H}}\leq C$ almost surely for $b=1,\ldots,B$.
+Let $\hat{w}_{1},\ldots,\hat{w}_{B}\in\mathcal{H}$ be independent random variables in a Hilbert space $\mathcal{H}$ such that $\|\hat{w}_{b}-\mathbb{E}[\hat{w}_{b}]\|_{\mathcal{H}}\leq C$ almost surely for $b=1,\ldots,B$.
 Then for any $\delta>0$,
-$\mathbb{P}\left\{\bigg{\|}\frac{1}{B}\sum_{b=1}^{B}(\hat{w}_{b}-\mathbb{E}\hat{w}_{b})\bigg{\|}_{\mathcal{H}}\leq C\sqrt{\frac{1}{2B}\log\left(\frac{2e^{2}}{\delta}\right)}\right\}\geq 1-\delta.$
+$\mathbb{P}\left\{\bigg{\|}\frac{1}{B}\sum_{b=1}^{B}(\hat{w}_{b}-\mathbb{E}[\hat{w}_{b}])\bigg{\|}_{\mathcal{H}}\leq C\sqrt{\frac{1}{2B}\log\left(\frac{2e^{2}}{\delta}\right)}\right\}\geq 1-\delta.$
 In particular,
-$\mathbb{E}\left[\bigg{\|}\frac{1}{B}\sum_{b=1}^{B}(\hat{w}_{b}-\mathbb{E}\hat{w}_{b})\bigg{\|}_{\mathcal{H}}^{2}\right]\leq\frac{C^{2}e^{2}}{B}.$
+$\mathbb{E}\left[\bigg{\|}\frac{1}{B}\sum_{b=1}^{B}(\hat{w}_{b}-\mathbb{E}[\hat{w}_{b}])\bigg{\|}_{\mathcal{H}}^{2}\right]\leq\frac{C^{2}e^{2}}{B}.$
 Note that, aside from the additional factor $e^{2}$, this result is the same as what we would obtain for the case $\mathcal{H}=\mathbb{R}$.
@@
 for some $t=(t_{1},\dots,t_{K})\in\mathbb{R}^{K}$ with $\|t\|_{1}\leq R$.
-For instance, returning to our earlier example, if $\mathcal{W}=\Delta_{K-1}$ (and the norm is given by $\mathsf{d}_{\mathsf{TV}}$) then we can take $v_{j}$ to be the $j$th canonical basis vector, and then 2 holds, with this same value of $K$, and with $R=1$ and $\rho=0$.
 Here $R$ is playing a role analogous to the radius, $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})$, in the Hilbert space setting.
+For instance, returning to our earlier example, if $\mathcal{W}=\Delta_{K-1}$ (and the norm is given by $\mathsf{d}_{\mathsf{TV}}$) then we can take $v_{j}$ to be the $j$th canonical basis vector.
+Then 2 holds, with this same value of $K$, and with $R=1$ and $\rho=0$.
 Theorem 10.
@@
 Problem Setting 2(b): the continuous case
-Next, consider a continuous setting, where $\mathcal{A}$ returns a density on $[0,1]$, given by some function $\hat{w}:[0,1]\rightarrow[0,\infty)$ satisfying $\int_{t=0}^{1}\hat{w}(t)\;\mathsf{d}t=1$.
-In this case we have $\mathsf{d}_{\mathsf{TV}}(w,w^{\prime})=\frac{1}{2}\int_{t=0}^{1}|w(t)-w^{\prime}(t)|\;\mathsf{d}t$.
+Next, consider a continuous setting, where $\mathcal{A}$ returns a density on $[0,1]$, given by some function $\hat{w}:[0,1]\rightarrow[0,\infty)$ satisfying $\int_{t=0}^{1}\hat{w}(t)\;\textnormal{d}t=1$.
+In this case we have $\mathsf{d}_{\mathsf{TV}}(w,w^{\prime})=\frac{1}{2}\int_{t=0}^{1}|w(t)-w^{\prime}(t)|\;\textnormal{d}t$.
 To formalize this setting, consider the Banach space
-$\mathcal{B}=L_{1}([0,1]):=\left\{f:[0,1]\to\mathbb{R}:\int_{t=0}^{1}|f(t)|\;\mathsf{d}t<\infty\right\},$
-equipped with the norm $\|f\|=\frac{1}{2}\int_{t=0}^{1}|f(t)|\;\mathsf{d}t$ (note that this differs from the usual $\ell_{1}$ norm by the factor $\frac{1}{2}$, in order to agree with total variation distance).
+$\mathcal{B}=L_{1}([0,1]):=\left\{f:[0,1]\to\mathbb{R}:\int_{t=0}^{1}|f(t)|\;\textnormal{d}t<\infty\right\},$
+equipped with the norm $\|f\|=\frac{1}{2}\int_{t=0}^{1}|f(t)|\;\textnormal{d}t$ (note that this differs from the usual $\ell_{1}$ norm by the factor $\frac{1}{2}$, in order to agree with total variation distance).
 Suppose that our algorithm returns densities that are further constrained to be $L$-Lipschitz, so that we have
@@
 As for Problem Setting 2(a), we again begin by asking whether the Hilbert space result of Theorem 6 might already be sufficient in this setting.
-Indeed, while $\mathcal{B}=L_{1}([0,1])$ is not a Hilbert space, the Lipschitz constraint on $\mathcal{W}$ means that $\int_{t=0}^{1}|w(t)|^{2}\;\mathsf{d}t<\infty$ for all $w\in\mathcal{W}$, meaning that $\mathcal{W}$ can be viewed as a subset of the Hilbert space $L_{2}([0,1])$, equipped with norm $\|f\|_{\mathcal{H}}=\|f\|_{L_{2}}=(\int_{t=0}^{1}|f(t)|^{2}\;\mathsf{d}t)^{1/2}$.
+Indeed, while $\mathcal{B}=L_{1}([0,1])$ is not a Hilbert space, the Lipschitz constraint on $\mathcal{W}$ means that $\int_{t=0}^{1}|w(t)|^{2}\;\textnormal{d}t<\infty$ for all $w\in\mathcal{W}$, meaning that $\mathcal{W}$ can be viewed as a subset of the Hilbert space $L_{2}([0,1])$, equipped with norm $\|f\|_{\mathcal{H}}=\|f\|_{L_{2}}=(\int_{t=0}^{1}|f(t)|^{2}\;\textnormal{d}t)^{1/2}$.
 A straightforward calculation verifies that $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})\propto L^{1/4}$, meaning that by Theorem 6, derandomized bagging satisfies mean-square stability with respect to $\|\cdot\|_{L_{2}}$ for $\beta^{2}\propto L^{1/2}/n$.
@@
 Output space $\mathcal{W}$
-Stability guarantee via
-Theorem 6 (Hilbert space)
-Stability guarantee via
-Theorem 10 (Banach space)
+Stability guarantee for $\mathsf{d}_{\mathsf{TV}}$
+via Theorem 6
+(Hilbert space)
+Stability guarantee for $\mathsf{d}_{\mathsf{TV}}$
+via Theorem 10
+(Banach space)
 Problem Setting 2(a):
@@
 We apply sklearn.tree.DecisionTreeRegressor to train the regression trees, setting max_depth=50 and leaving all other parameters at their default values.
-We evaluate the stability of the learned function in the $L_{2}$ norm $\|\hat{w}\|_{L_{2}}=\int_{[0,1]^{d}}\hat{w}(x)\textnormal{d}x$.
+We evaluate the stability of the learned function in the $L_{2}$ norm $\|\hat{w}\|_{L_{2}}=(\int_{[0,1]^{d}}\hat{w}(x)^{2}\textnormal{d}x)^{1/2}$.
 Experiment 2 (Synthetic controls real data example)
@@
 When we run SCM on a subset of the $n=16$ countries, we assign a weight of 00 to each omitted country.
-We run SCM using the Python package pysyncon and make extensive use of their reproduction of the Abadie et al.
-(2015) study.
+We run SCM using the Python package pysyncon and make extensive use of their reproduction of the original study Abadie et al.
+(2015).
 We evaluate the stability of these weights in the Euclidean norm $\|\cdot\|_{2}$.
@@
 Higher curves thus represent greater instability.
-In the right panels we also plot the stability guarantee of Theorem 6 for Experiments 1-3 and Theorem 10 for Experiment 4 (the guarantees for derandomized bagging).
+In the right panels we also plot the stability guarantee of Theorem 6 for Experiments 1–3 and Theorem 10 for Experiment 4 (the guarantees for derandomized bagging).
 For all four experiments, $m=n/2$ and $B=10,000$.
@@
 In the right panels of Figure 1, for a fixed data set and algorithm, we measure stability by plotting, for each value of $\varepsilon$, the smallest value of $\delta$ such that the algorithm is $\left(\varepsilon,\delta\right)$-stable:
-$\delta=\frac{1}{n}\sum_{i=1}^{n}1_{\|\hat{w}-\hat{w}^{\setminus i}\|>\varepsilon}.$
+$\delta=\frac{1}{n}\sum_{i=1}^{n}1_{\|\hat{w}-\hat{w}^{\setminus i}\|\geq\varepsilon}.$
 Furthermore, in Table 2, we record the mean-square stability
@@
 Synthetic controls (SCM) $0.0532$ $0.0125$
-Spectral analysis (LSSA) $40.033$ $0.3017$
-Softmax $0.0744$ $0.0014$
+Spectral analysis (LSSA) $40.033$ $0.0793$
+Softmax $0.0186$ $0.0006$
 Table 2: Results for mean-square stability.
@@
 $\leq\frac{1}{1-p_{\mathcal{Q}_{n}}}\mathbb{E}\left[\left\|\sum_{i\in[n]}(\hat{w}-\hat{w}^{\setminus i})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right\|_{\mathcal{H}}^{2}\right]^{1/2}\cdot\mathbb{E}\left[\left\|\hat{w}^{r}-\hat{w}\right\|_{\mathcal{H}}^{2}\right]^{1/2}$
-$\leq\frac{\mathsf{rad}_{\mathcal{H}}(\mathcal{W})}{1-p_{\mathcal{Q}_{n}}}\mathbb{E}\left[\left\|\sum_{i\in[n]}(\hat{w}-\hat{w}^{\setminus i})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right\|_{\mathcal{H}}^{2}\right]^{1/2},$
-with the last step holding because we have
+$\leq\frac{\mathsf{rad}_{\mathcal{H}}(\mathcal{W})}{1-p_{\mathcal{Q}_{n}}}\mathbb{E}\left[\left\|\sum_{i\in[n]}(\hat{w}-\hat{w}^{\setminus i})\cdot\left(\mathbf{1}\{i\in r\}-p_{\mathcal{Q}_{n}}\right)\right\|_{\mathcal{H}}^{2}\right]^{1/2}.$
+The last step holds because we have
 $\mathbb{E}\left[\|\hat{w}^{r}-\hat{w}\|_{\mathcal{H}}^{2}\right]=\inf_{w\in\mathcal{W}}\mathbb{E}\left[\|\hat{w}^{r}-w\|_{\mathcal{H}}^{2}\right]\leq\inf_{w\in\mathcal{W}}\sup_{w^{\prime}\in\mathcal{W}}\|w^{\prime}-w\|_{\mathcal{H}}^{2}=\mathsf{rad}^{2}_{\mathcal{H}}(\mathcal{W}),$
-where the first equality holds since $\hat{w}=\mathbb{E}[\hat{w}^{r}]$.
+where the first equality holds since $\hat{w}=\mathbb{E}[\hat{w}^{r}]$ minimizes mean-square error.
 To calculate the last remaining expected value, we first note that since we have assumed symmetry and nonpositive covariance for $\mathcal{Q}_{n}$ (see 1), we have $\textnormal{Cov}\left(\mathbf{1}\{i\in r\},\mathbf{1}\{j\in r\}\right)=-q$ for all $i\neq j$, for some $q\geq 0$.
@@
 Then, solving the quadratic,
-$\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}\cdot A+\frac{A^{2}}{3}\ \Longrightarrow\ \sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq A^{2}\cdot\left(\frac{1+\sqrt{7/3}}{2}\right)^{2}.$
+$\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq\sqrt{\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}}\cdot A+\frac{A^{2}}{3}\ \Longrightarrow\ \sum_{i\in[n]}\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\leq A^{2}\cdot\left(\frac{1+\sqrt{7/3}}{2}\right)^{2}\leq 1.6A^{2}.$
 This completes the proof of the theorem. ∎
@@
 Let $\mathcal{B}=L_{1}(\mathbb{N})$ (equipped with the norm $\|x\|=\frac{1}{2}\sum_{i\geq 1}|x_{i}|$), and let $\mathcal{W}=\Delta_{\infty}$.
-Then there exists a (deterministic) algorithm $\mathcal{A}:\cup_{n\geq 0}\mathcal{Z}^{n}\times[0,1]\to\mathcal{W}$, such that for any distribution $\mathcal{Q}_{n}$ satisfying 1, derandomized bagging fails to satisfy mean-square stability with respect to $\mathsf{d}_{\mathsf{TV}}$ for any $\beta<p$.
+Then there exists a (deterministic) algorithm $\mathcal{A}:\cup_{n\geq 0}\mathcal{Z}^{n}\times[0,1]\to\mathcal{W}$, such that for all $n\geq 1$, for any distribution $\mathcal{Q}_{n}$ satisfying 1, derandomized bagging fails to satisfy mean-square stability with respect to $\mathsf{d}_{\mathsf{TV}}$ for any $\beta<p_{\mathcal{Q}_{n}}$.
 In other words, stability results of the type shown in Theorem 10 guarantee a bound on $\beta^{2}$ (the stability parameter), uniformly over all algorithms $\mathcal{A}$, that vanishes as $n\to\infty$.
@@
 $S(\mathcal{D})=\left\{i\in\mathbb{N}:\sum_{j\in[m]}\mathbf{1}\{Z_{j}=i\}\geq 1\right\},$
-the set of values appearing in the data set.
-Then define the algorithm as $\mathcal{A}(\mathcal{D};\xi)=\mathbf{e}_{\psi(S(\mathcal{D}))}\in\mathcal{W},$ where $\mathbf{e}_{j}$ is the $j$th canonical basis vector, with a $1$ in the $j$th location and 00’s elsewhere.
+the set of unique values appearing in the data set.
+Then define $\mathcal{A}(\mathcal{D};\xi)=\mathbf{e}_{\psi(S(\mathcal{D}))}\in\mathcal{W},$ where $\mathbf{e}_{j}$ is the $j$th canonical basis vector, with a $1$ in the $j$th location and 00’s elsewhere.
 Now consider data set $\mathcal{D}=(1,\dots,n)$.
@@
 $=p_{\mathcal{Q}_{n}},$
-since $\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{i\in(S(\mathcal{D}^{r}))\}=\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{i\in r\}=p_{\mathcal{Q}_{n}}$.
+since $\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{i\in S(\mathcal{D}^{r})\}=\mathbb{P}_{r\sim\mathcal{Q}_{n}}\{i\in r\}=p_{\mathcal{Q}_{n}}$.
 Since this holds for every $i$, we have $\frac{1}{n}\sum_{i\in[n]}\|\hat{w}-\hat{w}^{\backslash i}\|^{2}=p_{\mathcal{Q}_{n}}^{2}$, which completes the proof.
@@
 […]
-$\delta\epsilon^{2}\geq 1.6\left[R\sqrt{\frac{2(1+\eta_{n})\log(2K)}{n}\left(\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{4\eta_{n}}{(1-p_{\mathcal{Q}_{n}})^{2}}\right)}+R\sqrt{\frac{1}{n}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}+2\rho\cdot p_{\mathcal{Q}_{n}}\right]^{2}.$
+$\delta\varepsilon^{2}\geq 1.6\left[R\sqrt{\frac{2(1+\eta_{n})\log(2K)}{n}\left(\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{4\eta_{n}}{(1-p_{\mathcal{Q}_{n}})^{2}}\right)}+R\sqrt{\frac{1}{n}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}+2\rho\cdot p_{\mathcal{Q}_{n}}\right]^{2}.$
 Proof.
@@
 Corollary 15.
-Under the notation and assumptions of Theorem 10, let $B\geq 1$ be fixed.
-Then bagging, run with base algorithm $\mathcal{A}$ and resampling distribution $\mathcal{Q}_{n}$ and with $B$ bags, has mean-square stability $\beta^{2}$ with respect to $\|\cdot\|_{\mathcal{H}}$, with
-$\beta^{2}:=\inf_{\varepsilon>0}\Bigg{\{}(1+\epsilon)\cdot 1.6\bigg{[}R\sqrt{\frac{2(1+\eta_{n})\log(2K)}{n}\left(\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{4\eta_{n}}{(1-p_{\mathcal{Q}_{n}})^{2}}\right)}+R\sqrt{\frac{1}{n}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}\\ {}+2\rho\cdot p_{\mathcal{Q}_{n}}\bigg{]}^{2}{}+(1+\epsilon^{-1})\cdot\left(\frac{8KR^{2}}{B}+8\rho^{2}\right)\Bigg{\}}$
+[…]
+$\beta^{2}:=\inf_{\varepsilon>0}\Bigg{\{}(1+\varepsilon)\cdot 1.6\bigg{[}R\sqrt{\frac{2(1+\eta_{n})\log(2K)}{n}\left(\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}+\frac{4\eta_{n}}{(1-p_{\mathcal{Q}_{n}})^{2}}\right)}+R\sqrt{\frac{1}{n}\cdot\frac{p_{\mathcal{Q}_{n}}}{1-p_{\mathcal{Q}_{n}}}}\\ {}+2\rho\cdot p_{\mathcal{Q}_{n}}\bigg{]}^{2}{}+(1+\varepsilon^{-1})\cdot\left(\frac{8KR^{2}}{B}+8\rho^{2}\right)\Bigg{\}}$
 Proof.
@@
 We have
-$\frac{1}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\right]=\frac{1}{n}\sum_{i\in[n]}\mathbb{E}\left[\|(\hat{w}-\tilde{w})-(\hat{w}^{\setminus i}-\tilde{w}^{\setminus i})+(\tilde{w}-\tilde{w}^{\setminus i})\|^{2}\right]\\ \leq\frac{2(1+\epsilon^{-1})}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}-\tilde{w}\|^{2}\right]+\frac{2(1+\epsilon^{-1})}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}^{\setminus i}-\tilde{w}^{\setminus i}\|^{2}\right]+\frac{1+\epsilon}{n}\sum_{i\in[n]}\|\tilde{w}-\tilde{w}^{\setminus i}\|^{2},$
-since $(a+b+c)^{2}\leq(1+\epsilon)a^{2}+(1+\epsilon^{-1})(b+c)^{2}\leq(1+\epsilon)a^{2}+2(1+\epsilon^{-1})(b^{2}+c^{2})$, for any $a,b,c\geq 0$.
-The last term in the upper bound is bounded by Theorem 10.
-Now we bound the remaining terms.
+$\frac{1}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}-\hat{w}^{\setminus i}\|^{2}\right]=\frac{1}{n}\sum_{i\in[n]}\mathbb{E}\left[\|(\hat{w}-\tilde{w})-(\hat{w}^{\setminus i}-\tilde{w}^{\setminus i})+(\tilde{w}-\tilde{w}^{\setminus i})\|^{2}\right]\\ \leq\frac{1+\varepsilon}{n}\sum_{i\in[n]}\|\tilde{w}-\tilde{w}^{\setminus i}\|^{2}+\frac{2(1+\varepsilon^{-1})}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}-\tilde{w}\|^{2}\right]+\frac{2(1+\varepsilon^{-1})}{n}\sum_{i\in[n]}\mathbb{E}\left[\|\hat{w}^{\setminus i}-\tilde{w}^{\setminus i}\|^{2}\right],$
+[…]
 First, let $\hat{w}^{\prime}$ be an i.i.d. copy of $\hat{w}$. Under the same notation as in the proof of Theorem 10, let $\|\cdot\|_{*}$ denote the dual norm. Then